### v1

# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

Abstract
Many engineering and scientific workflows depend on expensive black-box evaluations, requiring decision-making that simultaneously improves performance and reduces uncertainty. Bayesian optimization (BO) and Bayesian experimental design (BED) offer powerful yet largely separate treatments of goal-seeking and information-seeking, providing limited guidance for hybrid settings where learning and optimization are intrinsically coupled. We propose pragmatic curiosity, a hybrid learning-optimization paradigm derived from active inference, in which actions are selected by minimizing the expected free energy—a single objective that couples pragmatic utility with epistemic information gain. We demonstrate the practical effectiveness and flexibility of pragmatic curiosity on various real-world hybrid tasks, including constrained system identification, targeted active search, and composite optimization with unknown preferences. Across these benchmarks, pragmatic curiosity consistently outperforms strong BO-type and BED-type baselines, delivering higher estimation accuracy, better critical-region coverage, and improved final solution quality.
Active inference, Hybrid learning-optimization
1 Introduction
Engineering and scientific applications often rely on expensive black-box evaluations to identify high-performing designs or desirable system states. Bayesian optimization (BO) accelerates this process when the primary objective is to reach a specified goal (Shahriari et al., 2016; Frazier, 2018), whereas Bayesian experimental design (BED) prioritizes information acquisition about unknown system parameters (Rainforth et al., 2023). Both methodologies leverage probabilistic models and acquisition criteria that quantify the utility of evaluating unknown configurations, tailored to either optimization or learning objectives. Despite their individual successes and the explosion of research in each field, their disconnection creates a vacuum for a broad class of hybrid problems that routinely require simultaneously both seeking knowledge and achieving goals.
For many real-world applications, such as goal-directed planning (Lookman et al., 2019), environmental monitoring (Konakovic Lukovic et al., 2020), and targeted material design (Matsumoto et al., 2025), learning and optimization are not separate phases but deeply intertwined objectives. This challenge fundamentally arises across tasks with increasing complexity in terms of both epistemic consideration (i.e., from parametric to non-parametric models) and pragmatic evaluation (i.e., from known to unknown goals):
(1) Constrained System Identification, where the epistemic desire of precisely learning a system’s parameters is governed by the pragmatic need to keep experiments within safe or valid operational bounds (e.g., avoiding sensor saturation, or dangerous chemical reaction). This type of task can be found in numerous applications, including environmental monitoring (Konakovic Lukovic et al., 2020) and catalyst design (Zhong et al., 2020).
(2) Targeted Active Search, where the pragmatic objective to discover regions that meet certain criteria (e.g., system failure modes, or specific performance ranges) requires epistemic curiosity to explore the region’s shape, size, and boundaries. Example applications can be found in failure discovery (Ramanagopal et al., 2018) and medical monitoring (Malkomes et al., 2021).
(3) Composite Bayesian Optimization, where the pragmatic goal is to find an optimal design based on a user’s hidden preferences—a task that is impossible without first being epistemically curious about the user’s objectives themselves. Such scenarios commonly arise in simulation-based design (González and Zavala, 2025; Coelho et al., 2025) and A/B testing (Bakshy et al., 2018)
Classically, to address these hybrid problems, practitioners have been forced to choose between specialized tools, and accommodate problem-specific adaptation by leveraging information-gain criteria to enhance optimization and vice versa. On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems. Hvarfner et al. (2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the model hyperparameter even as it searches for the optimum. On the BED (also known as Bayesian active learning, BAL) side, Smith et al. (2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution. These methods highlight growing synergy between optimization and learning, but they remain task-specific, and rarely generalize across categories.
In this paper, we propose pragmatic curiosity: a hybrid learning-optimization paradigm derived from active inference (AIF) (Friston, 2010; Friston et al., 2017). AIF prescribes action selection by minimizing expected free energy (EFE), a single objective that combines (i) a pragmatic term that favors preferred outcomes and (ii) an epistemic term that favors information gain. We prove that EFE minimization provides a unified perspective of various acquisition strategies: by specifying preferences, observation models, and approximations, the resulting criterion recovers BO-like and BED-like behaviors as limiting regimes.
Under this paradigm, seeking knowledge and achieving goals are not treated as competing objectives, but as two inseparable facets of a single imperative to minimize EFE. These two drives are balanced by a coefficient called curiosity, which sets the trade-off between learning and optimization. The formal role of curiosity in guaranteeing both self-consistent learning (i.e., posterior convergences to the truth) and no-regret optimization (i.e., with bounded cumulative regret) is provided theoretically in Li et al. (2026).
This paper, instead, demonstrates the practical effectiveness and flexibility of this paradigm in handling a broad class of complex hybrid problems often ignored by standard methods, including tasks with evolving goals (where conditions change over time) and implicit goals (where objectives are not defined a priori). We conduct experiments structured around the three problem classes mentioned above, drawing on applications that include environmental monitoring in plume fields (Konakovic Lukovic et al., 2020), failure detection in autonomous driving scenarios (Ramanagopal et al., 2018), and distributed energy resource allocation in power grids (Kianmehr et al., 2019).
The empirical results reveal a consistent pattern of superior performance, demonstrating our framework’s advantages in its ability to solve complex, hybrid objectives. In constrained system identification tasks, our algorithm achieved near-perfect estimation accuracy while requiring up to 40% fewer queries than other methods. For targeted active search tasks, it demonstrated a more effective exploration strategy, discovering a crucial 10% more of the critical failure region within the same budget. Most notably, in tasks with unknown user preferences, our approach always successfully learned the underlying objective where other baselines fail to capture. Together, these findings validate the power of our unified approach, showing that a principled balance between pragmatic and epistemic drives leads to tangible gains across diverse and challenging problem settings.
In summary, our main contributions are the following:
  • A unified perspective of various acquisition strategies through the lens of active inference.
  • A pragmatic curiosity paradigm for general hybrid learning-optimization problems.
  • A comprehensive empirical validation across three canonical real-world problem categories with diverse hybrid learning-optimization objectives.
2 Preliminaries
2.1 Bayesian Optimization
Given an unknown objective function $f:\mathcal{X}\mapsto\Re$, BO seeks to identify the input $x^{\ast}$ that maximizes the objective $f$ over an admissible set of queries $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x).$ To achieve this goal, BO relies on a surrogate model that provides a probabilistic representation of the objective $f$, and uses this information to compute an acquisition function to drive the selection of the most promising sample to query.
Surrogate model. We assume the available information regarding the objective function $f$ be stored in the dataset $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, where $y_{t}\sim\mathcal{N}(f(x_{t}),\sigma^{2}(x_{t}))$ is the noisy observation of the objective function by assuming the noise follows a zero-mean normal distribution with a standard deviation $\sigma$. The surrogate model depicts possible explanations of $f$ as $f(x)\sim p(f(x)|\mathcal{D}_{t})$ applying a joint distribution over its behavior at each sample $x\in\mathcal{X}$. In Bayesian inference, the prior distribution of the objective $p(f(x))$ is combined with the likelihood function $p(\mathcal{D}_{t}|f(x))$ to compute the posterior distribution $p(f(x)|\mathcal{D}_{t})\propto p(\mathcal{D}_{t}|f(x))p(f(x))$, representing the updated beliefs about $f(x)$. Typically, Gaussian processes (GPs) have been widely used as the surrogate model for BO due to their efficient posterior sampling that enables cheap, gradient-based optimization of the acquisition function to propose new query points. GP is specified by a joint normal distribution $p(f(x)|\mathcal{D}_{t})=\mathcal{N}(\mu_{t}(x),\kappa_{t}(x,x^{\prime}))$ with mean $\mu_{t}(x)$ and kernel function $\kappa_{t}(x,x^{\prime})$, where $\mu_{t}(x)$ represents the prediction and $\kappa_{t}(x,x^{\prime})$ the associated uncertainty.
Acquisition function. The surrogate model is utilized to decide the next sample $x_{t+1}\in\mathcal{X}$ through the maximization of an acquisition function $\alpha:\mathcal{X}\mapsto\Re$, i.e., $x_{t+1}=\arg\max_{x\in\mathcal{X}}\alpha(x|\mathcal{D}_{t})$, where $\alpha(x|\mathcal{D}_{t})$ provides a measure of the improvement that the next query is likely to provide with respect to (w.r.t.) the current surrogate model of the objective function. Many acquisition functions have been proposed, including probability of improvement (Močkus, 1975), expected improvement (Jones et al., 1998), upper confidence bound, and various entropy search methods (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Hvarfner et al., 2022b; Neiswanger et al., 2021), as well as practical approaches to optimize them (Wilson et al., 2018).
2.2 Bayesian Experimental Design
Rather than optimizing an objective function $f(x)$, the purpose of BED is to sequentially select a set of experimental designs $x\in\mathcal{X}$ and gather outcomes $y$, to maximize the amount of information obtained about certain parameters of interest, denoted as $\theta$. The parameters $\theta$ can correspond to some explicit model parameters, or any implicitly defined quantity of interest (e.g., the optimum of a function, the output of an algorithm, or future downstream predictions).
Based on the current history of experiments $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, BED seeks to find the next experimental design $x_{t+1}$ by maximizing the expected information gain (EIG) (Chaloner and Verdinelli, 1995) that a potential experimental outcome $y_{t+1}$ can provide about $\theta$, measured in terms of expected entropy reduction of the posterior distribution of $\theta$:
$\begin{split}\text{EIG}(x|\mathcal{D}_{t})&=H[p(\theta|\mathcal{D}_{t})]-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[H[p(\theta|\mathcal{D}_{t}\cup(x,y))]]\\ &=I(\theta;(x,y)|\mathcal{D}_{t}),\end{split}$
where $H(\cdot)$ and $I(\cdot)$ denote the entropy and mutual information, respectively.
3 A Unified View of Acquisition Strategies
The acquisition strategies in BO typically lead to goal-directed behavior, where the (implicit) goal is the optimum of a certain (unknown) objective function. On the contrary, the acquisition strategies in BED encourage information-seeking behavior, aiming to gather the maximum amount of information about certain parameters of interest. Although they both can be viewed as realizations of adaptive sampling (Di Fiore et al., 2023), there are no transferable approaches between these two domains due to distinct directives (Hvarfner et al., 2025).
In this section, we show that these two seemingly competing imperatives can be naturally balanced through the principles of active inference (AIF).
3.1 Active Inference as Expected Free Energy Minimization
We specify a probabilistic surrogate model $q(\cdot)$ to capture the relationship between an outcome $y$ and a decision variable $x$ based on a set of parameters $s$ that are of interest, which factorizes as
$q(x,y,s)\coloneq p(x,y|s)q(s),$
where $q(x,y,s)$ is a joint probability distribution over $(x,y,s)$ based on a surrogate distribution $q(s)$ over $s$. We use $q(\cdot)$ to explicitly distinguish this model to be an “surrogate” of the “true” model $p(\cdot)$.
One special aspect of AIF is the way it formalizes goals. Instead of specifying the goal with additional variables related to “rewards” or “costs”, AIF directly encodes the preferences over possible outcomes $y$ through a probability distribution $p(y)$. In this distribution, outcomes with higher probabilities are treated as more rewarding. The deviation between observed outcomes and those desired is measured through an information-theoretic quantity known as self-information or surprisal: $\log p(y)$. Intuitively, it’s a measure of how unexpected an outcome $y$ is, given a prior preference distribution $p(y)$. Consistent with the intuitive notion of surprise, lower probability events generate higher surprisal values. Then the surprisal incurred from an observed outcome $y$ is quantified by
$\begin{split}-\log p(y)&=-\log\int p(y,s)ds\\ &=-\log\int\frac{p(y,s)q(s)}{q(s)}ds\\ &=-\log\mathbb{E}_{q(s)}\left[\frac{p(y,s)}{q(s)}\right]\\ &\leq-\mathbb{E}_{q(s)}\left[\log\frac{p(y,s)}{q(s)}\right]=F,\end{split}$ (1)
where the last inequality follows Jensen’s inequality, which states that the expectation of a logarithm is always less than or equal to the logarithm of an expectation.
The right-hand side of (1) is called the variational free energy (VFE), whose name arises from the fact that $F$ resembles the Helmholtz free energy in physics. We can see that VFE is always greater than or equal to surprisal (i.e., it is an upper bound on surprisal). In machine learning, the sign of VFE is typically reversed, making it an evidence lower bound (ELBO). Maximizing the ELBO is a commonly used optimization approach in machine learning (Titsias, 2009).
To formulate a strategy for decision-making, we need to consider the decision variable $x$ and the outcomes that result from the choice of actions. Since the future outcomes have not yet occurred, we resort to examining the expectation of surprisal over predicted outcomes based on a predictive distribution $q(y|x)$:
$\begin{split}-\mathbb{E}_{q(y|x)}&\log p(y|x)\\ \leq-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(y,s|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(s|x,y)p(y|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y,s|x)}\log p(y|x)\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y|x)}\log p(y|x)=G,\end{split}$ (2)
where the right-hand side of (2) is denoted as the expected free energy (EFE).
Theorem 3.1.
When using a surrogate model $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is considered as the true model constructed from all available data $\mathcal{D}_{t}$, $G$ can be simplified as
$\begin{split}G&=-\mathbb{E}_{q(y|x)}\left[D_{\text{KL}}(q(s|x,y)||q(s))\right]-\mathbb{E}_{q(y|x)}\log p(y)\\ &=\underbrace{-I_{q}(s;(x,y))}_{\text{epistemic}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y)}_{\text{pragmatic}},\end{split}$ (3)
where $I_{q}(\cdot)$ represents the mutual information given the surrogate model $q(\cdot)$.
Proof.
See Appendix B. ∎
We can see that by construction, EFE strikes a balance between information-seeking and goal-directed behavior under some prior preferences. It bounds the difference between epistemic value (about parameters) and pragmatic value (about outcomes), which captures the imperative to maximize the epistemic value (i.e., information gain about latent states), from interactions with the environment, while maximizing the pragmatic value (i.e., expected preference alignment), regarding prior preferences. This crucial aspect of AIF effectively addresses the “explore-exploit dilemma” because the imperatives for exploration and exploitation are just two aspects of EFE:
Pragmatic Value (Exploitation): This term encourages goal-directed behavior by favoring actions expected to yield preferred outcomes. Encoded by a prior distribution over desired observations, it functions similarly to a utility or reward function in reinforcement learning (Millidge et al., 2020), driving the agent to exploit its current knowledge to achieve its goals.
Epistemic Value (Exploration): This term promotes information-seeking behavior by favoring actions expected to reduce uncertainty about the underlying system maximally. It quantifies the expected information gain about the model’s parameters, driving the agent to explore the environment to refine its world model.
3.2 Reinterpretation of Acquisition Strategies in BO and BED
Crucially, minimizing EFE serves as a unifying umbrella principle. Many classic acquisition strategies in BO and BED can be reinterpreted as special cases of minimizing EFE, as shown in Table 1.
Table 1: A unified view of different acquisition strategies in BO and BED, where $x^{\ast}$, $y^{\ast}$ represent the true optimal solution and value, respectively, and $\hat{y}$ is the best value observed in $\mathcal{D}_{t}$.
Acquisition Strategy Acquisition Function $\alpha(x)$ Expected Free Energy
Parameters $s$ Preferences $p(y)$
Expected Information Gain (Chaloner and Verdinelli, 1995) $I(\theta;(x,y)|\mathcal{D}_{t})$ $\theta$ -
Entropy Search (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014) $I(x^{*};(x,y)|\mathcal{D}_{t})$ $x^{*}$ -
Max-value Entropy Search (Wang and Jegelka, 2017) $I(y^{*};(x,y)|\mathcal{D}_{t})$ $y^{*}$ -
Joint Entropy Search (Hvarfner et al., 2022b) $I((x^{*},y^{*});(x,y)|\mathcal{D}_{t})$ $(x^{*},y^{*})$ -
Bayesian Algorithm Execution (Neiswanger et al., 2021) $I(\mathcal{O}_{\mathcal{A}}(f);(x,y)|\mathcal{D}_{t})$ $\mathcal{O}_{\mathcal{A}}(f)$ -
GP-Upper Confidence Bound $\mu_{t}(x)+\beta^{1/2}\sigma_{t}(x)$ $f_{\mathcal{X}}$ $\exp\{y\}$
[…]
The reinterpretations of most of the acquisition strategies in Table 1 are straightforward according to their definitions. However, placing a rather intuitive GP-UCB strategy within this framework seems implicit. To reveal their connection, we rely on the following lemma:
Lemma 3.2.
Let $\mathbf{X}\subseteq\mathcal{X}$ be a subset of inputs, and the corresponding function values evaluated at those inputs be denoted as $f_{\mathbf{X}}$. Given a historical dataset $\mathcal{D}_{t}$, and new measurements $\mathbf{Y}$ observed at $\mathbf{X}$, the mutual information
$I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D}_{t})=I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}_{t}),$
for any (finite or infinite) set $\mathcal{X}$.
Proof.
See Appendix C. ∎
Then if we assume constant Gaussian noises $\mathcal{N}\sim(0,\sigma^{2})$ for the observations, we have
$\begin{split}I(f_{x};y|\mathcal{D}_{t})&=H(y|\mathcal{D}_{t})-H(y|f_{x},\mathcal{D}_{t})\\ &=\frac{1}{2}\log(1+\sigma^{-2}\sigma_{t}^{2}(x)),\end{split}$
where $\sigma_{t}^{2}(x)$ is the variance evaluated on the GP model $p(f_{x}|\mathcal{D}_{t})$.
When further assuming that the GP kernel $\kappa_{t}(x,x^{\prime})\leq 1,\forall x,x^{\prime}\in\mathcal{X}$, then $0\leq\sigma_{t}^{2}(x)\leq\kappa_{t}(x,x^{\prime})\leq 1$, which gives
$\log(1+\sigma^{-2}\sigma_{t}^{2}(x))\geq\log(1+\sigma^{-2})\sigma_{t}^{2}(x).$
If we choose $\beta=\frac{1}{2}\log(1+\sigma^{-2})$, then the epistemic term in EFE, i.e., $I(f_{x};y|\mathcal{D}_{t})$, provides an upper bound of the square of the exploration term $\beta^{1/2}\sigma_{t}(x)$ in GP-UCB.
This reveals the close relationship between GP-UCB and AIF, showing that even a seemingly pure intuitive strategy, like GP-UCB, can have rather rigorous mathematical foundations underlying it.
4 A New Paradigm to Derive Acquisition Functions
The preference distribution $p(y)$ in Table 1 can be interpreted as a softmax transformation of a time-varying value function over the outcomes. Generalizing from this observation, we propose a new paradigm for deriving acquisition functions applicable to a wider class of hybrid problems other than traditional BO and BED.
Grounded on Boltzmann distribution (also called Gibbs distribution) in statistical mechanics, we define an Boltzmann operator $\mathcal{B}_{\beta}$ that maps an energy function $h(z)$ over domain $\mathcal{Z}$ into a Boltzmann distribution:
$(\mathcal{B}_{\beta}[h])(z)\coloneq\frac{e^{-h(z)/\beta}}{\int_{\mathcal{Z}}e^{-h(z)/\beta}dz},$
where $\beta$ is called a temperature parameter (in allusion to statistical mechanics). A higher temperature results in a more uniform output distribution (i.e.“more random” with higher entropy), while a lower temperature results in a sharper output distribution, with the maximizers of $h(z)$ dominating.
Then, for any pre-defined energy function $h(y|\mathcal{D}_{t})$, we can transfer it into a Boltzmann distribution and use it as the preference distribution $p(y)$ in (3):
$\begin{split}G&=-I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\log{e^{-h(y|\mathcal{D}_{t})/\beta}}]+Z\\ &=-I(s;(x,y)|\mathcal{D}_{t})+\frac{1}{\beta}\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]+Z,\end{split}$
where $p(y|x,\mathcal{D}_{t})$ is the predictive distribution of a surrogate model constructed from the historical data $\mathcal{D}_{t}$, and $Z=\log\int_{\mathcal{Y}}e^{-h(y|\mathcal{D}_{t})/\beta}dy$ is a normalization constant independent of $y$.
Therefore, we propose a new acquisition paradigm, pragmatic curiosity, by minimizing the EFE, i.e.,
$x_{t+1}=\arg\max_{x\in\mathcal{X}}\{\beta_{t}I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]\},$
where $\beta_{t}\geq 0$ is a parameter that modulates the degree of curiosity, and $h(y|\mathcal{D}_{t})$ is a problem-dependent energy function that captures the goal (in terms of regret w.r.t. a certain outcome $y$).
The curiosity $\beta_{t}$ plays an important role in balancing the performance between learning and optimization. Li et al. (2026) derive a formal theoretical guarantee of posterior consistency and bounded cumulative regret under a sufficient curiosity condition, and provide practical design guidelines for choosing $\beta_{t}$ in different settings.
This paper instead examines how a problem-dependent energy function, $h(y|\mathcal{D}_{t})$, provides a flexible mechanism for explicitly representing uncertainty over the goals themselves. This added expressiveness enables acquisition design for a broad class of complex settings that standard methods often overlook, including tasks with evolving goals (where conditions change over time) and implicit goals (not specified a priori). The effectiveness and versatility of this formulation on such tasks are demonstrated in the next section.
5 Experiments
In this section, we exemplify the benefits and variability of the proposed acquisition strategy by performing experiments across three categories that differ substantially from one another. These problems originate from distinct literature within BO and BED, and none of them are canonical BO or BED tasks. Thus each of them is evaluated against a different set of BO-type (optimization-focused) and BED-type (learning-focused) baselines appropriate for that specific task, ensuring a fair and rigorous comparison.
Figure 1: Performance evaluation for constrained system identification on environmental monitoring in 2d plume fields. Error bars represent $\pm 1$ std over 5 seeds.
5.1 Parametric Models with Known Invariant Goals
We first consider a simplest task where the goals are known and invariant, and we have a decent understanding of the model, where only certain finite parameters are unknown.
One typical problem structure is the constrained system identification whose objective is to precisely learn unknown system parameters, $\theta$, when valid observations can only be gathered under specific operational constraints, quantified by $C(y)\leq 0$. For instance, when monitoring a chemical plume, the goal is to identify its source ($\theta$) while ensuring that sensor measurements ($y$) do not exceed a saturation threshold ($C(y)$).
In such cases, it is straightforward to choose the interested state as $s=\theta$, and the energy function as $h(y|\mathcal{D}_{t})=\mathbb{I}(C(y)>0)$. This leads to an acquisition function:
$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(\theta;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}\left[\mathbb{I}(C(y)>0)\right].\end{split}$
Tasks. We perform experiments on a real-world environmental monitoring problem in 2d plume fields where the sensors have a saturation threshold $y_{max}$ (i.e., $C(y)=y-y_{max}$) (Detailed settings and hyper-parameter choices in Appendix D.2). We consider three types of monitoring tasks: (a) locating the unknown source location; (b) estimating unknown wind direction and strength; and (c) identifying the active sources in fields with multiple sources.
Baselines. We compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by choosing the point that leads to the least probability of violating the constraint (BO-type); and (c) EIG about the unknown parameter (BED-type).
Evaluations. We evaluate the performance from both epistemic and pragmatic perspectives: (a) estimation accuracy; and (b) constraint violations.
Results. Fig. 1 shows that our approach achieves consistently stronger query efficiency over baselines while respecting all operational constraints, with cumulative constraint violation always being zero. This advantage is especially evident in the source localization task, where the drive for information and the need to satisfy constraints create opposing pressures. Resolving this conflict, our approach reaches near-perfect estimation using up to 40% fewer queries than competing methods.
Figure 2: Performance evaluation for targeted active search on failure discovery in autonomous driving scenarios. Error bars represent $\pm 1$ std over 4 seeds.
5.2 Non-Parametric Models with Known Evolving Goals
Next, we consider a more challenging setting where the task condition evolves, and the model is fully black-box such that we need to resort to a non-parametric model (e.g., GPs).
One such example is the targeted active search in multi-objective design problems. The objectives are viewed as metrics where specific ranges carry particular significance, and the goal is to design experiments that maximize coverage of these important regions $\mathcal{S}$.
Adopted from Malkomes et al. (2021), we assume that the experimental design problems possess a sense of known resolution, such as simulation accuracy or manufacturing precision/tolerance, and any outcome within distance $\delta$ of another does not convey extra information about $\mathcal{S}$. Similar to Malkomes et al. (2021), we define the coverage neighborhood of any $y$ as $\mathbb{C}_{\delta}(y):=\{y^{\prime}:d(y,y^{\prime})<\delta\}$. And the coverage neighborhood of a set of points $Y$ is defined as $\mathbb{C}_{\delta}(Y):=\bigcup_{y\in Y}\mathbb{C}_{\delta}(y)$.
Since the goal is to cover as much volume of the interested region $\mathcal{S}$ as possible, $h(y|\mathcal{D}_{t})$ can be chosen as $h(y|\mathcal{D}_{t})=\text{Vol}(\mathcal{S})-\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})$, where $Y$ contains all the history of outcomes. Similarly, coverage over the design configurations can be flexibly incorporated by expanding the interested state into the whole input space, i.e., $s=f_{\mathcal{X}}$. With non-parametric models like GPs, we can leverage Lemma 3.2 and derive an acquisition function:
$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(f_{x};y|\mathcal{D}_{t})+\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})].\end{split}$
Tasks. We perform experiments on a real-world failure discovery problem in autonomous driving scenarios where the perception module, a YOLO detector, may fail due to multiple causes, which could potentially lead to collisions (3d input-2d output). We consider three target sets with decreasing volume, i.e., Target Set 1 $\supset$ Target Set 2 $\supset$ Target Set 3 (Detailed settings and hyper-parameter choices in Appendix D.3).
Baselines. We again compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by maximizing coverage volume in metrics space (BO-type); and (c) EIG by maximizing coverage volume in parameter space (BED-type).
Evaluations. We evaluate the performance by inspecting the coverage volume over both spaces: (a) metrics space ($\mathbb{C}_{\delta}(Y)$); and (b) parameter space ($\mathbb{C}_{\delta}(X)$).
Results. As shown in Fig. 2, our AIF algorithm effectively balances coverage of both the parameter and metrics spaces. This capability is especially impactful for smaller target sets where the search is more difficult. In the most challenging case (Target Set 3), our method identifies nearly 10% more of the critical failure region compared to the leading baseline.
Figure 3: Performance evaluation for composite BO with unknown preferences. Error bars represent $\pm 1$ std over 20 seeds for vehicle safety, penicillin, and 5 for energy resource allocation.
5.3 Non-Parametric Models with Unknown Goals
Finally, we look into the most difficult setting where both the models and goals are black-box.
One practical scenario arises from the composite Bayesian optimization in multi-objective optimization problems. The objectives are weighted by a preference function $g(y)$ that is unknown a priori and must be simultaneously learned during the optimization process.
As assumed in Lin et al. (2022), we can estimate the unknown preference function $g(y)$ by asking a decision-maker (DM) to express preferences over pairs of outcomes $(y_{1},y_{2})$. Let $z(y_{1},y_{2})\in\{1,2\}$ indicate whether the DM preferred the first or second outcome offered. Following Chu and Ghahramani (2005), we assume that the DM’s responses are distributed according to a probit likelihood: $L(z(y_{1},y_{2})=1|g(y_{1}),g(y_{2}))=\Phi(\frac{g(y_{1})-g(y_{2})}{\sqrt{2}\lambda}),$ where $\lambda$ is a hyperparameter, and $\Phi$ is the standard normal CDF.
Taking the joint exploration and exploitation of the preference function $g(y)$ into consideration, we can choose $h(y|\mathcal{D}_{t})$ as the EFE of the preference model, i.e., $h(y|\mathcal{D}_{t})=\text{EFE}(g(y))$, which naturally extends the acquisition function into a nested structure:
[…]
where $x=[x_{1},x_{2}]$ are two jointly evaluated candidates, and $\gamma,\beta\geq 0$.
Tasks. We perform experiments on three real-world problems, including vehicle safety (5d-3d), penicillin production simulator (7d-3d), and distributed energy resource allocation in power grids (40d-4d) (Detailed settings and hyper-parameter choices in Appendix D.4).
Baselines. Different from BOPE in Lin et al. (2022), which separates and iterates the preference exploration and experimentation stages, our approach unifies the exploration and exploitation of both outcome and preference models simultaneously. In its best setting (i.e., switch stages and update models at each step), their acquisition strategy reduces to a special case of our proposed (4) with $\beta=\gamma=0$. Thus, we design a set of baselines to investigate the effect that each component in (4) plays for the joint exploration and exploitation of hierarchical models. We ablate the full AIF strategy (g + g IG + f IG) by (a) removing the last term (EIG); (b) removing the first term (g + g IG); and (c) removing the first two terms (g only).
Evaluations. We evaluate their performance by inspecting the best preference of all collected outcomes using the true preference function $g(y)$.
Results. Fig. 3 illustrates our AIF approach’s outstanding ability to learn unknown preference functions, a key advantage over baselines that often fail due to ill-directed queries. As the tasks become more complex and noisy (from (a) to (c)), our results demonstrate that every component of our acquisition function (4) plays an irreplaceable role in achieving optimal performance. This benefit is most stark in the energy resource allocation task, where competing methods failed to capture any meaningful preference model while our approach consistently succeeded.
Benefits of joint learning and optimization. To highlight the benefits of jointly learning and optimizing, rather than separating these into stages, we compare our method against several BOPE variants from Lin et al. (2022) that use different stage-wise design choices. Detailed experimental setups, design choices, plots, and analyses are provided in Appendix D.4.3. The results show that our method naturally balances exploration and exploitation at each step and consistently discovers higher-preference regions, whereas the BOPE variants are highly sensitive to how the stages are configured. Consequently, stage-wise approaches like BOPE require careful manual tuning of these choices, while our unified formulation automates this trade-off and is therefore more amenable to higher-order hierarchical models.
6 Conclusion and Limitation
We presented pragmatic curiosity, an AIF-based paradigm for hybrid learning-optimization under expensive black-box evaluations. By selecting actions via EFE minimization, the approach unifies goal-seeking and information-seeking within a single acquisition objective. Across constrained system identification, targeted active search, and composite optimization with unknown preferences, pragmatic curiosity consistently outperformed strong BO-type and BED-type baselines, improving estimation accuracy, critical-region coverage, and final solution quality under fixed budgets.
The limitation of our approach stems from the specification of the problem-dependent energy/preference model; misspecification can bias acquisition and degrade performance. Performance also inherits assumptions of the underlying surrogate and observation models; strong model mismatch or non-stationarity can compromise both uncertainty quantification and preference-induced guidance. Future work includes extending the evaluation of this paradigm to multi-agent, long-horizon, or multi-fidelity settings.
Impact Statement
This work advances sample-efficient decision-making for engineering and scientific systems where evaluations are expensive and objectives are coupled to uncertainty. By unifying goal-seeking and information-seeking within a single EFE objective, pragmatic curiosity provides a general acquisition-design template that can reduce experimental cost, accelerate discovery, and improve reliability in workflows such as safe system identification, targeted failure-region discovery, and preference-driven design. In domains like environmental monitoring, autonomous-system testing, and energy-resource allocation, the practical benefit is fewer costly or risky evaluations for a given level of performance and confidence, which can translate into reduced resource consumption and safer experimentation.
Potential negative impacts stem from misuse or mis-specification of the preference/energy model: poorly chosen goals or biased preferences can steer exploration toward harmful regions (e.g., unsafe operating conditions) or encode inequitable outcomes in preference-driven optimization. In addition, the ability to more efficiently identify system vulnerabilities could be misused in adversarial settings. These risks motivate careful deployment practices, including explicit safety constraints, transparent preference specification, robustness checks under model mismatch, and domain-appropriate oversight when the method is applied to safety-critical or socially sensitive applications.
Appendix A Related Work
Several independent attempts have been made to handle the hybrid objectives of goal seeking and learning. Below, we discuss some key prior works, originating from three viewpoints: (1) BO-based settings, (2) BED (also known as Bayesian active learning, BAL) based works, and (3) some works that have acknowledged the synergy and attempted to present a unifying viewpoint, using adaptive sampling and AIF-based approaches.
On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems by minimizing the ratio between squared expected single-period regret and a measure of information gain. Hvarfner et al. (2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the Gaussian process hyperparameter even as it searches for the optimum. Tu et al. (2022); Hvarfner et al. (2022a) propose an information-theoretic acquisition function for multi-objective BO, which considers the joint information gain for the optimal set of inputs and outputs. Overall, these works acknowledge and demonstrate the value of information gain in enhancing the capabilities of goal-seeking strategies, with an ultimate focus on optimization. However, they do not explicitly comment on the synergy of goal seeking and learning, or formalize the acquisition strategies as a way to explicitly balance both objectives.
On the BED (also known as Bayesian active learning, BAL) side, Smith et al. (2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution. Lookman et al. (2019) discusses the performance of various BAL and BO strategies for material design application, noting their connection with adaptive sampling, and simultaneous exploration and exploitation. This shows that the realistic application of BO/BED-based approaches often necessitates a formalized viewpoint unifying each distinct objective.
These efforts begin to show the growing synergy between BO and BED techniques, where information-gain criteria are leveraged to enhance optimization and vice versa. In search of a unifying perspective for these techniques, Di Fiore et al. (2023) views BO and BAL as symbiotic adaptive sampling techniques driven by common principles of ‘goal-driven learning’. However, their viewpoint is limited to BO and BED as separate realizations of adaptive sampling, and they do not discuss a way to combine both objectives in a single acquisition strategy.
In this paper, we present a unified perspective that arises naturally from the principles of Active Inference (AIF) (Friston, 2010; Friston et al., 2017). Originally developed in computational neuroscience, AIF has been successfully applied in fields such as robotics (Lanillos et al., 2021), cognitive science (Smith et al., 2022), ecosystem modeling (Friston et al., 2024), and goal-directed planning in partially observable environments (Matsumoto et al., 2025). These works leverage the ability of AIF to accommodate goal-directed and epistemic considerations, which we utilize as the backbone for our unifying mechanism discussed in this paper.
Appendix B Proof of Theorem 3.1
Proof.
Recall that
$G=\underbrace{-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]}_{\text{Term 1}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y|x)}_{\text{Term 2}}$
Since our purpose is to align the true outcome distribution given the decision $x$ (i.e., $p(y|x)$) with the prior preference distribution $p(y)$, we replace the second term of $G$ by $\mathbb{E}_{q(y|x)}\log p(y)$. It is commonly referred to as pragmatic value, which scores the anticipated surprisal of the future outcomes given a predictive distribution $q(y|x)$.
For the first term of $G$, add and subtract $\log q(s|x,y)$ inside the expectation, we get
$\begin{split}\text{Term 1}&=-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)\right]+\mathbb{E}_{q(s|x,y)}\left[\log q(s|x,y)-\log q(s|x)\right]\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\underbrace{-D_{\text{KL}}(q(s|x,y)||p(s|x,y))}_{\text{Term 3}}+\underbrace{D_{\text{KL}}(q(s|x,y)||q(s|x))}_{\text{Term 4}}\right].\end{split}$ (5)
Since the surrogate model $q(s)$ usually will not be updated until a new outcome $y$ is observed, it doesn’t depend on the decision variable $x$ solely, i.e., $q(s|x)=q(s)$. Thus Term 4 becomes $D_{\text{KL}}(q(s|x,y)||q(s))$, which is commonly referred to as the epistemic value, or the expected information gain of a state when it is conditioned on expected observations.
Term 3 is, in general, intractable since it depends on the true posterior distribution $p(s|x,y)$ that is often unknown. The best surrogate model we can have is $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is constructed from all available data $\mathcal{D}_{t}$. Thus if we treat this surrogate model $q(s)$ as the best approximation of the true model $p(s)$, i.e., $q(s)\approx p(s)$, and the surrogate and true models both follow the Bayes’ rule, i.e., $q(s|x,y)=\frac{p(x,y|s)q(s)}{\int p(x,y|s)q(s)ds},\quad p(s|x,y)=\frac{p(x,y|s)p(s)}{\int p(x,y|s)p(s)ds},$ then Term 3 vanishes. ∎
Appendix C Proof of Lemma 3.2
To prove Lemma 3.2, we first introduce a lemma:
Lemma C.1.
For any (finite or infinite) set $\mathcal{X}$, after a few new measurements $(\mathbf{X},\mathbf{Y}),\mathbf{X}\subseteq\mathcal{X}$ the KL divergence between $p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))$ and $p(f_{\mathcal{X}}|\mathcal{D})$ is
$D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]=\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],$
where $L(\mathbf{Y}|f_{\mathbf{X}})$ is the likelihood of observations, $L(\mathbf{Y})=\int L(\mathbf{Y}|f_{\mathbf{X}})p(f_{\mathbf{X}}|\mathcal{D})df_{\mathbf{X}}$ the marginal likelihood.
Proof.
We first consider the case where $\mathcal{X}$ is finite to provide an intuitive insight. The KL divergence between $p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))$ and $p(f_{\mathcal{X}}|\mathcal{D})$ can be given as
$\begin{split}&D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]\\ =&D_{\text{KL}}[p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})]\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\int\frac{p(f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (6)
where $L(\mathbf{Y}|f_{\mathbf{X}})$ is the likelihood of observations, $L(\mathbf{Y})=\int L(\mathbf{Y}|f_{\mathbf{X}})p(f_{\mathbf{X}}|\mathcal{D})df_{\mathbf{X}}$ the marginal likelihood.
Now we move to the more general cases where $\mathcal{X}$ is infinite. In such cases, there is no useful infinite-dimensional Lebesgue measure with respect to an “infinite-dimensional vector” $f_{\mathcal{X}}$. Thus, we need to resort to a more general definition for KL divergence based on the Radon-Nikodym derivative:
Definition C.2.
If $P$ and $Q$ are probability measures over a set $\mathcal{X}$, and $P$ is absolutely continuous with respect to $Q$, then the KL divergence from $P$ to $Q$ is defined as
$D_{\text{KL}}[P||Q]=\int_{\mathcal{X}}\log(\frac{dP}{dQ})dP,$
where $\frac{dP}{dQ}$ is the Radon–Nikodym derivative of $P$ with respect to $Q$, and provided the expression on the right-hand side exists.
According to the measure-theoretic definition of Bayes’ theorem for a dominated model (Schervish, 1995), the Radon-Nikodym derivative of the posterior $P(\cdot):=p(\cdot|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))$ with respect to the prior $\hat{P}(\cdot):=p(\cdot|\mathcal{D)}$ is given as
$\frac{dP}{d\hat{P}}(f_{\mathcal{X}})=\frac{L(\mathbf{Y}|f_{\mathcal{X}})}{L(\mathbf{Y})}.$
Since the dataset $\mathbf{Y}$ is finite, so similar to previous, we restrict the likelihood to only depend on the finite dataset:
$\frac{dP}{d\hat{P}}(f_{\mathcal{X}})=\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}.$
Now the KL divergence between $\hat{P}$ and $P$ is quantified as
$\begin{split}&D_{\text{KL}}[P(f_{\mathcal{X}})||\hat{P}(f_{\mathcal{X}})]\\ =&\int_{f_{\mathcal{X}}}\log(\frac{dP}{d\hat{P}}(f_{\mathcal{X}}))dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathcal{X}})\\ =&\int_{f_{\mathbf{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathbf{X}})\\ =&\mathbb{E}_{\hat{P}(f_{\mathbf{X}})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (7)
which has the exact same form as (6).
Therefore, we can conclude that regardless of the set $\mathcal{X}$ being finite or infinite, the KL divergence between the prior and posterior only depends on the evaluations of the observed data. That is to say, whilst we are in fact quantifying the KL divergence between the full distributions, we only need to keep track of the distribution over finite function values $f_{\mathbf{X}}$. ∎
Now we are ready to prove Lemma 3.2.
Proof.
According to Lemma C.1,
$\begin{split}I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D})=&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}[D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]]\\ =&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{L(\mathbf{Y})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\mathbb{E}_{L(\mathbf{Y})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]L(\mathbf{Y})d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int L(\mathbf{Y}|f_{\mathbf{X}})\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[D_{\text{KL}}[L(\mathbf{Y}|f_{\mathbf{X}})||L(\mathbf{Y})]]\\ =&I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}).\end{split}$ (8)
Appendix D Experimental Details
This appendix provides a comprehensive overview of the simulation environment, model parameters, and hyper-parameter choices used to generate the results in this paper. The experiments were designed to be reproducible given the configurations outlined below.
D.1 Simulation Environment
All simulations for the perception failure evaluation in CARLA (Section D.3) were conducted on a Linux workstation with Ubuntu 22.04 LTS equipped with an Intel 13th Gen Core i7-13700KF CPU (16 cores, 24 threads, up to 5.4 GHz) and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). The system ran CARLA simulations using CUDA 12.2 and NVIDIA driver version 535.230.02.
All other experiments were run on a MacBook Pro equipped with an Apple M2 Pro processor (10-core CPU, 16-core GPU) and a 3024 × 1964 Retina display. The GPU supports Metal 3, and the system was used as-is without external accelerators.
All experiments were conducted using Python 3.9. The core scientific computing libraries utilized were:
  • BoTorch (Balandat et al., 2020)
  • GPyTorch (Gardner et al., 2018)
In all experiments, we utilized the built-in Monte Carlo sampler in Botorch for the optimization of acquisition functions. The Monte Carlo samples are drawn from the posteriors for each model to approximate the expectations of acquisition functions.
D.2 Constrained System Identification
D.2.1 Plume Field Model and Parameters
We consider the monitoring of a chemical plume field, where multiple plume sources generate plume particles that can be measured by sensors. The field function is represented by the rate of hits, defined as the average number of particles per unit time measured by the sensor at a certain location.
The rate of hits for a chemical plume source is given as:
[…]
The measurement $y$, i.e., the number of particles measured, is modeled as a Poisson random variable with $R_{\theta}(x)\Delta t$ as the rate parameter, which leads to a likelihood model as
$L_{\theta}(y|x)=\frac{\exp(-R_{\theta}(x)\Delta t)(R_{\theta}(x)\Delta t)^{y}}{y!},$
where $\Delta t$ is the time taken to obtain a measurement.
General Field and Sensor Parameters:
  • Grid Size: 100 × 100 units.
  • Sensor Size ($a$): 1.0 units.
  • Measurement Time ($\Delta t$): 1.0 seconds.
Source Parameters: Eight potential plume sources were defined for the experiments. Their specific physical parameters are detailed as:
Source $\theta$ $R_{s}$ $\gamma$ $V$ $D$
1 [20.0, 20.0] 100.0 50.0 [0.5, 0.5] 10
2 [30.0, 80.0] 100.0 60.0 [-0.3, 0.2] 15
3 [45.0, 55.0] 15.0 50.0 [0.5, 0.5] 10
4 [50.0, 50.0] 18.0 30.0 [-0.3, 0.2] 15
5 [55.0, 45.0] 16.0 40.0 [0.2, -0.4] 12
6 [48.0, 52.0] 17.0 35.0 [0.1, 0.1] 11
7 [52.0, 55.0] 14.0 45.0 [-0.1, -0.1] 13
8 [52.0, 52.0] 18.0 40.0 [0.1, -0.1] 13
D.2.2 Task-Specific Configurations
Source Localization.
  • Goal: Estimate the unknown location $\theta=[x,y]$ of a single active source.
  • Ground Truth: Source 1 was configured as the single active source.
  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
  • Hypothesis Space: A discrete grid of potential locations spanning [0,100]×[0,100] with a resolution of 5 units, resulting in a 20 × 20 grid of 400 hypotheses.
  • Degree of curiosity: $\beta=0.5$.
Wind Estimation.
  • Goal: Estimate the unknown wind vector $V=[v_{x},v_{y}]$ for a single source with a known location.
  • Ground Truth: Source 2 was used, with its true wind vector set to [-0.3, 0.2].
  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
  • Hypothesis Space: A discrete grid of potential wind vectors spanning [-1,1]×[-1,1] with a resolution of 0.1 units, resulting in a 20 × 20 grid of 400 hypotheses.
  • Degree of curiosity: $\beta=1.0$.
Active Source Identification.
  • Goal: Identify the subset of active sources from a set of six potential sources with known locations.
  • Ground Truth: Sources 3-8 were used, with Sources 3, 5, 6, and 8 were set as active.
  • Maximum Sensor Threshold ($y_{max}$): 30.0 hits/second.
  • Hypothesis Space: The set of all possible on/off combinations for the six sources. This is a discrete space with $2^{6}=64$ unique hypotheses.
  • Degree of curiosity: $\beta=5.0$.
D.3 Constrained Active Search
D.3.1 Perception in self driving simulation CARLA
We consider the failure discovery for YOLO-based object detection (Jiang et al., 2022; Redmon and Farhadi, 2018) in the CARLA simulator (Dosovitskiy et al., 2017).
This requires the generation of various scenarios in the environment using CARLA simulator. The environment is a composition of a static context and scenario $\phi$. We use the probabilistic programming framework Scenic (Fremont et al., 2019) for sampling scenarios with varying contextual information for a fixed scenario variable $\phi$. For the simulations presented in this paper, we used a publically available, pre-existing environment (Dreossi et al., 2019), which consists of the ego vehicle maneuvering on the road with two non-ego agents– two non-ego cars and a pedestrian crossing the road. We use YOLO object detection model to detect all non-ego agents in a scene. The generated scenario is seeded for reproducibility, so that for a given scenario $\phi$, the environment can be treated as a deterministic quantity. Each scenario is defined using $\phi=[b_{e},b_{l},s]$, where $b_{e},b_{l}\in[5,15]$ represent the braking threshold of the ego car and lead car (non-ego car in-front of ego car) (m) and $s\in[0,\pi/2]$ denotes the sun altitude angle (rad). Each of these quantities is normalized to be within $[0,1]$, and the normalized scenario is chosen as the decision variable $x$ for active inference.
We are interested in environment variables (scenarios) that lead to two specific types of failures– failure to detect non-ego agents due to large distance from ego vehicle (Failure 1), and failure to detect non-ego agents due to poor scene lighting (Failure 2). Fig. 4 shows examples of the discussed object detection failures we aim to discover.
Figure 4: Examples of missed object detection by YOLO due to two reasons considered in perception failure case study in Section D.3. Fig (left to right): example of Failure-1 (distance), Failure-2 (poor light) and Failure-1 and Failure-2 both in one scene (distance and poor light), respectively. Bounding boxes for detected objects (misdetections) shown in yellow (red) with detection confidence numbers. Each scene has two cars and a pedestrian.
D.3.2 LLM-based evaluation for CARLA
We use LLM to perform evaluations for whether a generated scene corresponds to a failure due to specific type.
Each evaluation for a specific scenario corresponds to $T=60$ steps of simulations. Images recorded from the camera view are used for object detection and classification at every $10$ steps using YOLO-v3 (Redmon and Farhadi, 2018), and the classified image are used as inputs to GiT (Wang et al., 2022) to predict a likely failure type based on fine-tuning data.
Results obtained from YOLO along-with the reports from GiT are used as inputs to GPT3 model for failure evaluation, which is queried $6$ times per evaluation, and assigns binary scores pertaining to each failure mode for each scene (camera image). The average value reported across 6 scenes is used to construct $c_{i}:\mathcal{Z}\to\mathbb{R}$ for $i=1,2$ as: $c_{i}(z)=\frac{1}{T}\sum_{t=1}^{T}b^{i}_{t}.$ Here $b^{i}_{t}\in\{0,1\}$ is a scene-specific binary evaluation provided by the LLM based on report generated by GiT to assess if an object detection failure is observed in a given scene and corresponds to Failure-$i$.
We use GPT-3.5 Turbo model for LLM-based binary evaluations, with each evaluation we query the LLM 6 times and combine the binary evaluations for all 6 runs. Note that usage of LLM is not a core part of our methodology and is only used to as a subjective evaluator.
We show the prompt used for failure evaluation of each scene using the LLM in the box below. Information shown in red and blue is obtained from GiT captioning system, and CARLA respectively. The output of the LLM is used to obtain a binary number for each scene which is composed to give a scenario specific cost function $c_{1},c_{2}$.
D.3.3 Task-Specific Configurations
  • Goal: We consider two cost functions $c_{1},c_{2}$ associated with each type of failure. The goal is to sample from the set $\Omega=\{z|c_{1}(z)\geq C_{1},c_{2}(z)\geq C_{2}\}$, and we consider three target sets defined by $C_{1}=C_{2}=0.1$,
    $C_{1}=C_{2}=0.5$, and
    $C_{1}=C_{2}=0.8$.
  • Degree of curiosity: $\beta=20.0$.
Prompt Used for CARLA evaluation You will be provides the analysis of YOLO Object detection on an image that was taken from the camera feed of CARLA simulator. The simulator is simulating a pedestrian crossing the road before a car in front of the ego car. There are two cars and one pedestrian in each image. The information provided: 1. Objects detected: List of objects detected by YOLO in the image. This list should have atleast one object from the Objects to detect list 2. Objects to detect list: One object with one of the following labels: ‘car’,‘truck’, One object with one of the following labels: ‘car’, ‘truck’,‘bus’,‘motorcycle’,‘bicycle’ One object with one of the following labels: ‘person’ 3. Reason: The reason is a brief explanation of the failure to detect all objects, if that happens, and is generated by a pre-trained GiT model in the form of captions for the image. We are looking to discover images where YOLO fails to detect an object due to bad light and/or large distance. If the list of objects detected has an object missing from the objects to detect list, look at the reason. The reason can have other components as well, but it can ‘only’ be considered as bad light if at least one of the objects was failed to detect strictly due to bad light. Similarly, the reason can have other components as well, but it can ‘only’ be considered as large distance if the reason contains the phrase ‘far away’. Follow the response instructions while responding. Response Instructions: Respond should be an integer 0, 1, 2, 3 or 4: 0 indicating that at least one object was missing from the ’objects to detect’ list, but the reason provided does not correspond to bad light or large distance. 1 indicating that an object was not detected and the reason provided corresponds to bad light only. 2 indicating that an object was not detected, and the reason corresponds to large distance only. 3 indicating that an object was not detected, and the reason corresponds to both large distance and bad light. 4 indicating all objects are detected. Do not provide explanation. Response format: Response: [integer], where integer = 0,1,2,3,4. The list of objects detected and reason for incomplete detection for the image are as follows: Objects detected: {objects} Reason: {reason}
D.4 Composite Bayesian Optimization
D.4.1 Preference Evaluation
We simulate human-in-the-loop or policy-driven decision-making via pairwise preference queries. That is, for selected pairs of outcomes $(y_{1},y_{2})$, a preference function indicates which design is preferred. These preferences are generated based on a latent utility function, not revealed to the optimizer.
An initial set of 1 pairwise preferences is randomly sampled to initialize the model. Each step of the optimization selects new pairs to query, guided by the used acquisition strategy.
D.4.2 Task-Specific Configurations
Vehicle Safety.
  • Goal: Optimize vehicle crash-worthiness.
  • Testbed: See Tanabe and Ishibuchi (2020) for details.
  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[1864.7202,11.8199,0.2904]$.
  • Degree of curiosity: $\beta=\gamma=1.0$.
Penicillin.
  • Goal: Maximize the penicillin yield while minimizing time to ferment and the CO2 byproduct.
  • Testbed: See Liang and Lai (2021) for details.
  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[25.935,57.612,935.5]$.
  • Degree of curiosity: $\beta=\gamma=1.0$.
Energy Resource Allocation.
  • Goal: Identify deployment strategies for Distributed Energy Resources (DERs) in Optimal Power Flow (OPF) that align with implicit ethical preferences across multiple performance dimensions detailed in the following table.
  • Testbed: IEEE 30-bus network in pandapower library.
  • Ground Truth Preference: $g(y)=a^{\intercal}y$, where a=[1,1,2,1]a=[1,-1,2,1]
  • Degree of curiosity: $\beta=\gamma=1.0$.
Performance Metrics Definition
Voltage Fairness Measures the variance in bus voltages across the network; lower variance implies more equitable voltage delivery.
Total Cost Combines capital expenditures for DER installation and operational costs related to reactive power support.
Priority Area Coverage Quantifies the share of power delivered to high-priority buses, such as rural or underserved regions.
Resilience Assesses the percentage of time that all bus voltages remain within safe operating limits under perturbations (e.g., load uncertainty or line outages).
D.4.3 Extended Baseline Comparison with BOPE
To highlight the benefits of jointly learning and optimizing, rather than separating these into stages, we extend the baseline comparison with BOPE from Lin et al. (2022).
The original BOPE framework is intentionally flexible and leaves many problem-specific design choices open, especially regarding how and when to switch between preference exploration and experimentation. In our comparison, we consider four representative stage-wise variants:
  • BOPE-I: A two-phase strategy that starts with qEUBO (preference-focused exploration) and switches to qNEI in the second half (objective-driven refinement), illustrating the effect of premature exploitation.
  • BOPE-II: A two-phase strategy that starts with qNEI (exploring the objective space) and switches to qEUBO in the second half (exploiting the learned preference model), using a frozen outcome snapshot for qEUBO.
  • BOPE-III: A qEUBO-only variant where experiments are selected by qEUBO with a newly sampled objective realization at each iteration, encouraging stronger exploration through objective variation.
  • BOPE-IV: A BOPE variant that uses standard preference exploration (qEUBO) and selects experiments exclusively with qNEI, fully refitting both outcome and preference GPs after each update.
Figure 5: Extended baseline comparison with BOPE for energy resource allocation. Error bars represent $\pm 1$ standard deviation over 5 seeds.
Figure 5 reports their preference scores over random sampling (RS). It is evident that our active-inference method for hierarchical GPs (HGP) consistently discovers higher-preference regions after a brief initial exploration phase, while BOPE variants are highly sensitive to their stage-wise design choices. BOPE-I, being preference-driven in the first phase, initially attains higher preference values but fails to balance exploration and exploitation, leading to poor convergence in the second half; its performance is also sensitive to the precise switching point. BOPE-II explores first and then optimizes, achieving better final performance than BOPE-I, but the strict separation between exploration and exploitation still yields suboptimal outcomes. BOPE-III mixes both aspects but remains more exploitation-centric, performing better than BOPE-I/II yet still below HGP. BOPE-IV and HGP share the idea of refitting both models at each step, but BOPE-IV converges to a lower-preference solution. In contrast, our acquisition strategy jointly leverages information from both the outcome and preference models at every iteration, leading to higher sample efficiency and more reliable discovery of high-preference regions.
Appendix E Use of Large Language Models
We use LLM to revise and polish drafts of the introduction, abstract, conclusion, and limitations sections to enhance clarity, flow, and accessibility for a broader audience.
We also use LLM as an evaluator for the experiments of failure discovery in CARLA, which is detailed in Appendix D.3.

### v2

@@
-Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference
+Pragmatic Curiosity: A Unified Framework for Hybrid Learning and Optimization via Active Inference
 Abstract
-Many engineering and scientific workflows depend on expensive black-box evaluations, requiring decision-making that simultaneously improves performance and reduces uncertainty.
-Bayesian optimization (BO) and Bayesian experimental design (BED) offer powerful yet largely separate treatments of goal-seeking and information-seeking, providing limited guidance for hybrid settings where learning and optimization are intrinsically coupled.
-We propose pragmatic curiosity, a hybrid learning-optimization paradigm derived from active inference, in which actions are selected by minimizing the expected free energy—a single objective that couples pragmatic utility with epistemic information gain.
-We demonstrate the practical effectiveness and flexibility of pragmatic curiosity on various real-world hybrid tasks, including constrained system identification, targeted active search, and composite optimization with unknown preferences.
-Across these benchmarks, pragmatic curiosity consistently outperforms strong BO-type and BED-type baselines, delivering higher estimation accuracy, better critical-region coverage, and improved final solution quality.
-Active inference, Hybrid learning-optimization
+Many engineering and scientific workflows rely on expensive black-box evaluations, requiring sequential decisions that must both improve task performance and reduce uncertainty.
+Bayesian optimization (BO) and Bayesian experimental design (BED) provide powerful but largely separate treatments of goal-directed optimization and information-seeking experimentation, leaving limited guidance for hybrid settings in which learning and optimization are intrinsically coupled.
+We propose Pragmatic Curiosity (PraC), a unified framework for hybrid learning and optimization via active inference.
+PraC evaluates candidate queries by trading information gain about a task-relevant latent symbol against an expected regret-based potential over outcomes.
+This formulation exposes three operational design choices: which latent quantity should be clarified, how task value is encoded as regret, and how strongly information gain should be exchanged against pragmatic value.
+We instantiate PraC across three regimes of increasing complexity: decision-oriented plume monitoring with fixed global symbols and known downstream losses, targeted active search with induced local symbols and evolving coverage goals, and composite Bayesian optimization with hierarchical regret learning under unknown preferences.
+Across these regimes, PraC reduces downstream decision risk, improves coverage of critical outcome regions, and jointly learns predictive and preference structures without relying on task-specific staging rules.
 1 Introduction
-Engineering and scientific applications often rely on expensive black-box evaluations to identify high-performing designs or desirable system states.
-Bayesian optimization (BO) accelerates this process when the primary objective is to reach a specified goal (Shahriari et al., 2016; Frazier, 2018), whereas Bayesian experimental design (BED) prioritizes information acquisition about unknown system parameters (Rainforth et al., 2023).
-Both methodologies leverage probabilistic models and acquisition criteria that quantify the utility of evaluating unknown configurations, tailored to either optimization or learning objectives.
-Despite their individual successes and the explosion of research in each field, their disconnection creates a vacuum for a broad class of hybrid problems that routinely require simultaneously both seeking knowledge and achieving goals.
-For many real-world applications, such as goal-directed planning (Lookman et al., 2019), environmental monitoring (Konakovic Lukovic et al., 2020), and targeted material design (Matsumoto et al., 2025), learning and optimization are not separate phases but deeply intertwined objectives.
-This challenge fundamentally arises across tasks with increasing complexity in terms of both epistemic consideration (i.e., from parametric to non-parametric models) and pragmatic evaluation (i.e., from known to unknown goals):
-(1) Constrained System Identification, where the epistemic desire of precisely learning a system’s parameters is governed by the pragmatic need to keep experiments within safe or valid operational bounds (e.g., avoiding sensor saturation, or dangerous chemical reaction).
-This type of task can be found in numerous applications, including environmental monitoring (Konakovic Lukovic et al., 2020) and catalyst design (Zhong et al., 2020).
-(2) Targeted Active Search, where the pragmatic objective to discover regions that meet certain criteria (e.g., system failure modes, or specific performance ranges) requires epistemic curiosity to explore the region’s shape, size, and boundaries.
-Example applications can be found in failure discovery (Ramanagopal et al., 2018) and medical monitoring (Malkomes et al., 2021).
-(3) Composite Bayesian Optimization, where the pragmatic goal is to find an optimal design based on a user’s hidden preferences—a task that is impossible without first being epistemically curious about the user’s objectives themselves.
-Such scenarios commonly arise in simulation-based design (González and Zavala, 2025; Coelho et al., 2025) and A/B testing (Bakshy et al., 2018)
-Classically, to address these hybrid problems, practitioners have been forced to choose between specialized tools, and accommodate problem-specific adaptation by leveraging information-gain criteria to enhance optimization and vice versa.
-On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems.
-Hvarfner et al.
-(2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the model hyperparameter even as it searches for the optimum.
-On the BED (also known as Bayesian active learning, BAL) side, Smith et al.
-(2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution.
-These methods highlight growing synergy between optimization and learning, but they remain task-specific, and rarely generalize across categories.
-In this paper, we propose pragmatic curiosity: a hybrid learning-optimization paradigm derived from active inference (AIF) (Friston, 2010; Friston et al., 2017).
-AIF prescribes action selection by minimizing expected free energy (EFE), a single objective that combines (i) a pragmatic term that favors preferred outcomes and (ii) an epistemic term that favors information gain.
-We prove that EFE minimization provides a unified perspective of various acquisition strategies: by specifying preferences, observation models, and approximations, the resulting criterion recovers BO-like and BED-like behaviors as limiting regimes.
-Under this paradigm, seeking knowledge and achieving goals are not treated as competing objectives, but as two inseparable facets of a single imperative to minimize EFE.
-These two drives are balanced by a coefficient called curiosity, which sets the trade-off between learning and optimization.
-The formal role of curiosity in guaranteeing both self-consistent learning (i.e., posterior convergences to the truth) and no-regret optimization (i.e., with bounded cumulative regret) is provided theoretically in Li et al.
-(2026).
-This paper, instead, demonstrates the practical effectiveness and flexibility of this paradigm in handling a broad class of complex hybrid problems often ignored by standard methods, including tasks with evolving goals (where conditions change over time) and implicit goals (where objectives are not defined a priori).
-We conduct experiments structured around the three problem classes mentioned above, drawing on applications that include environmental monitoring in plume fields (Konakovic Lukovic et al., 2020), failure detection in autonomous driving scenarios (Ramanagopal et al., 2018), and distributed energy resource allocation in power grids (Kianmehr et al., 2019).
-The empirical results reveal a consistent pattern of superior performance, demonstrating our framework’s advantages in its ability to solve complex, hybrid objectives.
-In constrained system identification tasks, our algorithm achieved near-perfect estimation accuracy while requiring up to 40% fewer queries than other methods.
-For targeted active search tasks, it demonstrated a more effective exploration strategy, discovering a crucial 10% more of the critical failure region within the same budget.
-Most notably, in tasks with unknown user preferences, our approach always successfully learned the underlying objective where other baselines fail to capture.
-Together, these findings validate the power of our unified approach, showing that a principled balance between pragmatic and epistemic drives leads to tangible gains across diverse and challenging problem settings.
-In summary, our main contributions are the following:
-  • A unified perspective of various acquisition strategies through the lens of active inference.
-  • A pragmatic curiosity paradigm for general hybrid learning-optimization problems.
-  • A comprehensive empirical validation across three canonical real-world problem categories with diverse hybrid learning-optimization objectives.
+Sequential decision-making under uncertainty often requires an agent to decide what to learn and what to do at the same time.
+Two classical paradigms have addressed these aspects from different directions.
+Bayesian optimization (BO) is primarily goal-directed: it selects queries in order to identify high-performing inputs of an unknown objective function (Močkus, 1975b; Jones et al., 1998a; Srinivas et al., 2012).
+Bayesian experimental design (BED) is primarily information-seeking: it selects experiments in order to reduce uncertainty about latent quantities of interest (Chaloner and Verdinelli, 1995a).
+Although both can be viewed as forms of adaptive sampling (Di Fiore et al., 2023), they are usually developed under different directives, with different acquisition rules and different criteria for success.
+This separation leaves limited guidance for problems in which the relevant uncertainty, the downstream objective, and the sampling policy must be designed jointly.
+In environmental monitoring (Konakovic Lukovic et al., 2020), a robot may need to learn the hidden state of a chemical plume while deciding where to dispatch a response team.
+In failure discovery (Ramanagopal et al., 2018; Parashar et al., 2025), an evaluator may need to explore uncertain scenarios while prioritizing outcomes that cover critical failure regions.
+In preference-guided design (González and Zavala, 2025; Coelho et al., 2025), an optimizer may need to learn not only how actions produce outcomes, but also how those outcomes should be valued.
+In all these settings, an agent must act before it fully knows, and it only learns by acting.
+The bottleneck is therefore not learning or optimization alone, but the meta-decision that connects them: deciding what knowledge is sufficient for action, and what actions are most informative for future decisions.
+Recent work has begun to address this coupling by choosing between specialized tools and accommodating problem-specific adaptation through leveraging information-gain criteria to enhance optimization and vice versa.
+Information-directed sampling (IDS) connects information and regret in online optimization and bandit settings (Russo and Van Roy, 2018).
+Self-correcting Bayesian optimization introduces a statistical distance-based active learning criterion into the BO loop to improve model learning during optimization (Hvarfner et al., 2023).
+Expected predictive information gain (EPIG) focuses active learning on predictive information that is relevant to a target distribution (Smith et al., 2023).
+Preference exploration for BO learns preference models over multi-objective outcomes, but typically relies on stage-wise choices between experimentation and preference learning (Jerry Lin et al., 2022).
+These methods highlight growing synergy between learning and optimization, but they often remain task-specific, and rarely generalize across categories.
+In this paper, we propose Pragmatic Curiosity (PraC), a unified framework for hybrid learning and optimization via active inference (AIF) (Friston, 2010; Friston et al., 2017).
+Originally developed in computational neuroscience, AIF prescribes action selection by minimizing expected free energy (EFE), a single objective that couples an epistemic drive for information gain with an pragmatic drive toward preferred outcomes.
+PraC translates this principle into a practical acquisition-design template: a candidate query is evaluated by both what it is expected to reveal about a task-relevant latent quantity and how its predicted outcome is expected to affect downstream action.
+In this sense, PraC formalizes pragmatic curiosity: curiosity is not the indiscriminate gathering of information, but the search for knowledge insofar as that knowledge can improve action.
+PraC not only serves as a unifying umbrella principle that re-interprets many classical acquisition rules in BO and BED, but also opens the door to practical acquisition design for a broader class of hybrid problems by exposing a triad of meta-decisions in such problems.
+First, the agent must decide what hidden structure is worth representing, since not all uncertainty is relevant to action.
+Second, it must decide how outcomes should be evaluated, since different downstream goals induce different notions of success, failure, risk, or shortfall.
+Third, it must decide how strongly unresolved uncertainty should influence the next action, since acting before knowing can be brittle while learning without regard to the goal can become aimless.
+Thus, hybrid decision-making is governed by a triad of meta-decisions in representation, evaluation, and regulation.
+In PraC, this triad is conceptualized as symbols, regrets, and curiosity.
+Symbols specify the latent quantities whose clarification can change what the agent ought to do.
+In parametric models, these may be fixed global symbols such as physical parameters or latent hypotheses; in non-parametric models, they may be induced local symbols tied to contemplated observations.
+Regret acts as a carrier of goal: it gives an operational representation of what matters by quantifying how an outcome falls short of what is desired.
+When the goal is known, a regret-based potential can be designed directly to encode domain knowledge; when the goal is unknown or implicit, the regret itself can be learned through a hierarchical structure of PraC.
+Curiosity specifies the exchange rate between clarifying knowledge and pursuing the current surrogate of the goal.
+Together, these three choices turn PraC from an abstract decision paradigm into a practical decision rule that accounts for diverse hybrid learning and optimization problems.
+This perspective also elevates curiosity from an ad-hoc exploration heuristic into an intrinsic regularizer that couples belief updating and decision-making.
+In many acquisition functions, the exploration weight appears as a fixed hyperparameter selected by empirical tuning.
+In PraC, curiosity has a semantic role: it determines how much pragmatic sacrifice is justified by epistemic clarification.
+This allows the effective epistemic pressure of the acquisition to be monitored and scheduled dynamically.
+We introduce a practical feedforward–feedback scheduler that calibrates curiosity from the relative scale of information gain and downstream pragmatic value, while suppressing curiosity as decision-relevant uncertainty decreases.
+In this way, curiosity becomes a fluid exchange rate that self-regulates as belief and action co-evolve.
+We instantiate PraC across three regimes of increasing complexity and empirically evaluate the role of each design axis.
+First, we study fixed global symbols with known downstream goals in decision-oriented monitoring.
+Here the latent symbol is a finite hypothesis, and the regret is the Bayes risk of a downstream response decision.
+PraC with dynamic curiosity scheduler achieves the lowest final Bayes risk in source response localization and active source prioritization, and obtains the best realized response loss in consequence-weighted dispatch, showing that curiosity can help resolve symbolic distinctions that matter for action rather than merely reducing uncertainty for its own sake.
+Second, we study induced local symbols with known evolving goals in targeted active search.
+Here the relevant symbol is induced by a non-parametric model, and the pragmatic value of a new outcome depends on what has already been covered.
+PraC coordinates exploration in the input space with coverage in the outcome space, and in the most challenging target set discovers nearly 10% more normalized target-region coverage than the strongest baseline (Malkomes et al., 2021a).
+This demonstrates the benefit of allowing symbols to be induced locally by contemplated observations rather than fixed globally in advance.
+Third, we study hierarchical symbols with unknown goals in composite Bayesian optimization with preferences.
+Here the agent must learn both how actions produce outcomes and how outcomes should be valued.
+PraC consistently outperforms ablated variants that remove components of the hierarchical acquisition, and further improves over BOPE-style stage-wise preference-exploration variants (Jerry Lin et al., 2022).
+These results show that a hierarchical regret design can jointly support outcome learning and preference learning without manually separating them into stages.
+Our contributions are as follows.
+  • We derive a practical active-inference acquisition template that couples information gain about a task-relevant latent symbol with an expected regret-based potential over outcomes.
+  • We identify three operational design axes—epistemic symbol, regret potential, and curiosity coefficient—and show how they recover, reinterpret, or approximate several BO/BED acquisition strategies while supporting new hybrid designs.
+  • We introduce a dynamic curiosity scheduler that calibrates the information–regret exchange rate from candidate-level acquisition statistics and posterior uncertainty.
+  • We demonstrate the framework across decision-oriented environmental monitoring, targeted active search, and preference-guided composite optimization, showing that the same principle adapts across parametric, non-parametric, and hierarchical regimes.
 2 Preliminaries
 2.1 Bayesian Optimization
-Given an unknown objective function $f:\mathcal{X}\mapsto\Re$, BO seeks to identify the input $x^{\ast}$ that maximizes the objective $f$ over an admissible set of queries $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x).$ To achieve this goal, BO relies on a surrogate model that provides a probabilistic representation of the objective $f$, and uses this information to compute an acquisition function to drive the selection of the most promising sample to query.
+Given an unknown objective function $f:\mathcal{X}\to\mathbb{R}$, BO seeks to identify the input $x^{\ast}$ that maximizes the objective $f$ over an admissible set of queries $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x).$ To achieve this goal, BO relies on a surrogate model that provides a probabilistic representation of the objective $f$, and uses this information to compute an acquisition function to drive the selection of the most promising sample to query.
 Surrogate model.
-We assume the available information regarding the objective function $f$ be stored in the dataset $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, where $y_{t}\sim\mathcal{N}(f(x_{t}),\sigma^{2}(x_{t}))$ is the noisy observation of the objective function by assuming the noise follows a zero-mean normal distribution with a standard deviation $\sigma$.
+We assume the available information regarding the objective function $f$ be stored in the dataset $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t})\}$, where $y_{t}\sim\mathcal{N}(f(x_{t}),\sigma^{2}(x_{t}))$ is the noisy observation of the objective function by assuming the noise follows a zero-mean normal distribution with a standard deviation $\sigma(x)$.
 The surrogate model depicts possible explanations of $f$ as $f(x)\sim p(f(x)|\mathcal{D}_{t})$ applying a joint distribution over its behavior at each sample $x\in\mathcal{X}$.
@@
 Acquisition function.
-The surrogate model is utilized to decide the next sample $x_{t+1}\in\mathcal{X}$ through the maximization of an acquisition function $\alpha:\mathcal{X}\mapsto\Re$, i.e., $x_{t+1}=\arg\max_{x\in\mathcal{X}}\alpha(x|\mathcal{D}_{t})$, where $\alpha(x|\mathcal{D}_{t})$ provides a measure of the improvement that the next query is likely to provide with respect to (w.r.t.) the current surrogate model of the objective function.
-Many acquisition functions have been proposed, including probability of improvement (Močkus, 1975), expected improvement (Jones et al., 1998), upper confidence bound, and various entropy search methods (Hennig and Christian J.
-Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Hvarfner et al., 2022b; Neiswanger et al., 2021), as well as practical approaches to optimize them (Wilson et al., 2018).
+The surrogate model is utilized to decide the next sample $x_{t+1}\in\mathcal{X}$ through the maximization of an acquisition function $\alpha:\mathcal{X}\to\mathbb{R}$, i.e., $x_{t+1}=\arg\max_{x\in\mathcal{X}}\alpha(x\mid\mathcal{D}_{t})$, where $\alpha(x\mid\mathcal{D}_{t})$ provides a measure of the improvement that the next query is likely to provide with respect to (w.r.t.) the current surrogate model of the objective function.
+Many acquisition functions have been proposed, including probability of improvement (Močkus, 1975a), expected improvement (Jones et al., 1998b), upper confidence bound, and various entropy search methods (Hennig and Christian J.
+Schuler, 2012; Hernández-Lobato et al., 2014; Wang and Jegelka, 2017; Hvarfner et al., 2022; Neiswanger et al., 2021), as well as practical approaches to optimize them (Wilson et al., 2018).
 2.2 Bayesian Experimental Design
@@
 The parameters $\theta$ can correspond to some explicit model parameters, or any implicitly defined quantity of interest (e.g., the optimum of a function, the output of an algorithm, or future downstream predictions).
-Based on the current history of experiments $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t}))\}$, BED seeks to find the next experimental design $x_{t+1}$ by maximizing the expected information gain (EIG) (Chaloner and Verdinelli, 1995) that a potential experimental outcome $y_{t+1}$ can provide about $\theta$, measured in terms of expected entropy reduction of the posterior distribution of $\theta$:
-$\begin{split}\text{EIG}(x|\mathcal{D}_{t})&=H[p(\theta|\mathcal{D}_{t})]-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[H[p(\theta|\mathcal{D}_{t}\cup(x,y))]]\\ &=I(\theta;(x,y)|\mathcal{D}_{t}),\end{split}$
+Based on the current history of experiments $\mathcal{D}_{t}:=\{(x_{1},y_{1}),\dots,(x_{t},y_{t})\}$, BED seeks to find the next experimental design $x_{t+1}$ by maximizing the expected information gain (EIG) (Chaloner and Verdinelli, 1995b) that a potential experimental outcome $y_{t+1}$ can provide about $\theta$, measured in terms of expected entropy reduction of the posterior distribution of $\theta$:
+$\begin{split}\text{EIG}(x\mid\mathcal{D}_{t})&=H[p(\theta|\mathcal{D}_{t})]-\mathbb{E}_{p(y\mid x,\mathcal{D}_{t})}[H[p(\theta|\mathcal{D}_{t}\cup(x,y))]]=I(\theta;(x,y)\mid\mathcal{D}_{t}),\end{split}$
 where $H(\cdot)$ and $I(\cdot)$ denote the entropy and mutual information, respectively.
-3 A Unified View of Acquisition Strategies
-The acquisition strategies in BO typically lead to goal-directed behavior, where the (implicit) goal is the optimum of a certain (unknown) objective function.
-On the contrary, the acquisition strategies in BED encourage information-seeking behavior, aiming to gather the maximum amount of information about certain parameters of interest.
-Although they both can be viewed as realizations of adaptive sampling (Di Fiore et al., 2023), there are no transferable approaches between these two domains due to distinct directives (Hvarfner et al., 2025).
-In this section, we show that these two seemingly competing imperatives can be naturally balanced through the principles of active inference (AIF).
+3 Pragmatic Curiosity: A Framework that Synthesizes Learning and Optimization
+Acquisition strategies in BO typically induce goal-directed behavior, where the implicit goal is to identify the optimum of an unknown objective function.
+By contrast, acquisition strategies in BED induce information-seeking behavior, aiming to collect the most informative data about latent quantities of interest.
+Although both can be viewed as instances of adaptive sampling (Di Fiore et al., 2023), they are usually developed under different directives and therefore admit few directly transferable design principles (Hvarfner et al., 2025).
+In this section, we synthesize these two seemingly competing imperatives within a unified framework through the lens of active inference (AIF).
 3.1 Active Inference as Expected Free Energy Minimization
-We specify a probabilistic surrogate model $q(\cdot)$ to capture the relationship between an outcome $y$ and a decision variable $x$ based on a set of parameters $s$ that are of interest, which factorizes as
-$q(x,y,s)\coloneq p(x,y|s)q(s),$
-where $q(x,y,s)$ is a joint probability distribution over $(x,y,s)$ based on a surrogate distribution $q(s)$ over $s$.
-We use $q(\cdot)$ to explicitly distinguish this model to be an “surrogate” of the “true” model $p(\cdot)$.
-One special aspect of AIF is the way it formalizes goals.
-Instead of specifying the goal with additional variables related to “rewards” or “costs”, AIF directly encodes the preferences over possible outcomes $y$ through a probability distribution $p(y)$.
-In this distribution, outcomes with higher probabilities are treated as more rewarding.
-The deviation between observed outcomes and those desired is measured through an information-theoretic quantity known as self-information or surprisal: $\log p(y)$.
-Intuitively, it’s a measure of how unexpected an outcome $y$ is, given a prior preference distribution $p(y)$.
-Consistent with the intuitive notion of surprise, lower probability events generate higher surprisal values.
-Then the surprisal incurred from an observed outcome $y$ is quantified by
-$\begin{split}-\log p(y)&=-\log\int p(y,s)ds\\ &=-\log\int\frac{p(y,s)q(s)}{q(s)}ds\\ &=-\log\mathbb{E}_{q(s)}\left[\frac{p(y,s)}{q(s)}\right]\\ &\leq-\mathbb{E}_{q(s)}\left[\log\frac{p(y,s)}{q(s)}\right]=F,\end{split}$ (1)
-where the last inequality follows Jensen’s inequality, which states that the expectation of a logarithm is always less than or equal to the logarithm of an expectation.
-The right-hand side of (1) is called the variational free energy (VFE), whose name arises from the fact that $F$ resembles the Helmholtz free energy in physics.
-We can see that VFE is always greater than or equal to surprisal (i.e., it is an upper bound on surprisal).
-In machine learning, the sign of VFE is typically reversed, making it an evidence lower bound (ELBO).
-Maximizing the ELBO is a commonly used optimization approach in machine learning (Titsias, 2009).
-To formulate a strategy for decision-making, we need to consider the decision variable $x$ and the outcomes that result from the choice of actions.
-Since the future outcomes have not yet occurred, we resort to examining the expectation of surprisal over predicted outcomes based on a predictive distribution $q(y|x)$:
-$\begin{split}-\mathbb{E}_{q(y|x)}&\log p(y|x)\\ \leq-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(y,s|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log\frac{p(s|x,y)p(y|x)}{q(s|x)}\right]\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y,s|x)}\log p(y|x)\\ =-&\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]\\ &-\mathbb{E}_{q(y|x)}\log p(y|x)=G,\end{split}$ (2)
-where the right-hand side of (2) is denoted as the expected free energy (EFE).
-Theorem 3.1.
-When using a surrogate model $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is considered as the true model constructed from all available data $\mathcal{D}_{t}$, $G$ can be simplified as
-$\begin{split}G&=-\mathbb{E}_{q(y|x)}\left[D_{\text{KL}}(q(s|x,y)||q(s))\right]-\mathbb{E}_{q(y|x)}\log p(y)\\ &=\underbrace{-I_{q}(s;(x,y))}_{\text{epistemic}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y)}_{\text{pragmatic}},\end{split}$ (3)
-where $I_{q}(\cdot)$ represents the mutual information given the surrogate model $q(\cdot)$.
-Proof.
-See Appendix B. ∎
-We can see that by construction, EFE strikes a balance between information-seeking and goal-directed behavior under some prior preferences.
-It bounds the difference between epistemic value (about parameters) and pragmatic value (about outcomes), which captures the imperative to maximize the epistemic value (i.e., information gain about latent states), from interactions with the environment, while maximizing the pragmatic value (i.e., expected preference alignment), regarding prior preferences.
-This crucial aspect of AIF effectively addresses the “explore-exploit dilemma” because the imperatives for exploration and exploitation are just two aspects of EFE:
-Pragmatic Value (Exploitation): This term encourages goal-directed behavior by favoring actions expected to yield preferred outcomes.
-Encoded by a prior distribution over desired observations, it functions similarly to a utility or reward function in reinforcement learning (Millidge et al., 2020), driving the agent to exploit its current knowledge to achieve its goals.
-Epistemic Value (Exploration): This term promotes information-seeking behavior by favoring actions expected to reduce uncertainty about the underlying system maximally.
-It quantifies the expected information gain about the model’s parameters, driving the agent to explore the environment to refine its world model.
-3.2 Reinterpretation of Acquisition Strategies in BO and BED
-Crucially, minimizing EFE serves as a unifying umbrella principle. Many classic acquisition strategies in BO and BED can be reinterpreted as special cases of minimizing EFE, as shown in Table 1.
+Consider a probabilistic surrogate model $q(\cdot)$ that captures the relationship between a decision variable $x$, an outcome $y$, and a collection of latent quantities of interest $s$, and factorizes as
+$q(x,y,s)\coloneq p(x,y\mid s)\,q(s),$
+where $q(s)$ is the surrogate belief over $s$. We use $q(\cdot)$ to distinguish the agent’s internal surrogate model from the underlying, generally inaccessible, true data-generating model $p(\cdot)$.
+In AIF, preferences over possible outcomes are encoded through a probability distribution $p(y)$.
+Outcomes assigned higher probability are treated as more preferred.
+The deviation between an observed outcome and those preferred by the agent is measured by its self-information, or surprisal, $\log p(y)$.
+Intuitively, surprisal quantifies how unexpected an outcome $y$ is under the preference distribution $p(y)$: less preferred outcomes incur larger surprisal.
+The surprisal associated with an outcome $y$ satisfies
+$\begin{split}-\log p(y)&=-\log\int p(y,s)\,ds=-\log\int\frac{p(y,s)q(s)}{q(s)}\,ds\\ &=-\log\mathbb{E}_{q(s)}\!\left[\frac{p(y,s)}{q(s)}\right]\leq-\mathbb{E}_{q(s)}\!\left[\log\frac{p(y,s)}{q(s)}\right]=F,\end{split}$ (1)
+where the inequality follows from Jensen’s inequality.
+The right-hand side of equation 1 is called the variational free energy (VFE), resembling the Helmholtz free energy in physics.
+It upper bounds surprisal and therefore provides a tractable surrogate objective for inference.
+In machine learning, the sign of VFE is often reversed, yielding the evidence lower bound (ELBO), whose maximization is a standard variational learning principle (Titsias, 2009).
+To obtain a decision rule, we must account not only for realized outcomes, but also for future outcomes that have not yet been observed.
+AIF does so by considering the expected surprisal of future outcomes under the predictive distribution $q(y\mid x)$:
+$\begin{split}-\mathbb{E}_{q(y\mid x)}&\log p(y\mid x)\leq-\mathbb{E}_{q(y,s\mid x)}\left[\log\frac{p(y,s\mid x)}{q(s\mid x)}\right]=-\mathbb{E}_{q(y,s\mid x)}\left[\log\frac{p(s\mid x,y)p(y\mid x)}{q(s\mid x)}\right]\\ =-&\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]-\mathbb{E}_{q(y,s\mid x)}\log p(y\mid x)\\ =-&\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]-\mathbb{E}_{q(y\mid x)}\log p(y\mid x)=G,\end{split}$ (2)
+where the right-hand side of equation 2 is denoted as the expected free energy (EFE). AIF prescribes action selection by minimizing this quantity.
+3.2 A Principled Decision Paradigm for Hybrid Learning and Optimization
+In its original form, the EFE in equation 2 is not yet an actionable acquisition rule, because it depends on the true posterior $p(s\mid x,y)$ and the true outcome model $p(y\mid x)$, both of which are generally unavailable a priori.
+We now turn it into a principled decision paradigm through two approximations that are natural from a decision-theoretic perspective.
+The EFE in equation 2 can be decomposed as
+$G=\underbrace{-\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x)\right]}_{\text{Term 1}}\;\underbrace{-\mathbb{E}_{q(y\mid x)}\log p(y\mid x)}_{\text{Term 2}}.$
+For Term 1, add and subtract $\log q(s\mid x,y)$ inside the expectation:
+$\begin{split}\text{Term 1}&=-\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)+\log q(s\mid x,y)-\log q(s\mid x)\right]\\ &=-\mathbb{E}_{q(y\mid x)}\mathbb{E}_{q(s\mid x,y)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)+\log q(s\mid x,y)-\log q(s\mid x)\right]\\ &=-\mathbb{E}_{q(y\mid x)}\left[\mathbb{E}_{q(s\mid x,y)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)\right]+\mathbb{E}_{q(s\mid x,y)}\left[\log q(s\mid x,y)-\log q(s\mid x)\right]\right]\\ &=-\mathbb{E}_{q(y\mid x)}\left[\underbrace{-D_{\text{KL}}(q(s\mid x,y)\|p(s\mid x,y))}_{\text{Term 3}}+\underbrace{D_{\text{KL}}(q(s\mid x,y)\|q(s\mid x))}_{\text{Term 4}}\right].\end{split}$
+Since the surrogate belief over $s$ is not updated until an observation is obtained, it does not depend on $x$ alone; hence $q(s\mid x)=q(s)$.
+Therefore, Term 4 becomes $D_{\text{KL}}(q(s\mid x,y)\|q(s))$, which is precisely the epistemic value: the information gained about the latent quantity $s$ from a prospective observation $(x,y)$.
+By contrast, Term 3 remains intractable because it depends on the true posterior $p(s\mid x,y)$.
+The best the agent can access is its current surrogate belief constructed from the historical data $\mathcal{D}_{t}$, namely $q(\cdot)=p(\cdot\mid\mathcal{D}_{t})$.
+This motivates the first approximation: we treat the surrogate belief as the best currently available approximation to the true model, i.e., $q(s)\approx p(s)$.
+If both the surrogate and true models update according to Bayes’ rule, $q(s\mid x,y)=\frac{p(x,y\mid s)q(s)}{\int p(x,y\mid s)q(s)ds},\quad p(s\mid x,y)=\frac{p(x,y\mid s)p(s)}{\int p(x,y\mid s)p(s)ds},$ then Term 3 vanishes under this approximation, and the resulting decision rule is Bayes-optimal with respect to the agent’s current belief.
+This is the only legitimate basis for action: a decision-maker cannot optimize with respect to an inaccessible ground truth, but only with respect to its present posterior approximation.
+Moreover, under standard consistency conditions, the discrepancy between surrogate and truth shrinks as more data are collected.
+The second obstacle lies in Term 2, which depends on the true predictive distribution $p(y\mid x)$.
+We therefore introduce a second approximation: we replace $p(y\mid x)$ by a preference distribution $p_{\text{pref}}(y)$ that does not depend on $x$.
+This converts Term 2 into $\mathbb{E}_{q(y\mid x)}\log p_{\text{pref}}(y)$, which evaluates preferred outcomes under the predictive model induced by choosing $x$.
+Intuitively, the agent now asks: if I take action $x$, what outcomes am I likely to see, and how desirable are those outcomes?
+To define a preference distribution from a task-level notion of desirability, we introduce a Boltzmann operator $\mathcal{B}_{\beta}$ that maps a nonnegative potential energy function $h:\mathcal{Z}\to\mathbb{R}_{\geq 0}$ into a probability distribution:
+$(\mathcal{B}_{\beta}[h])(z)\coloneq\frac{e^{-h(z)/\beta}}{\int_{\mathcal{Z}}e^{-h(z)/\beta}dz}.$
+This construction is inspired by the Boltzmann (or Gibbs) distribution in statistical mechanics.
+The parameter $\beta$ plays the role of temperature: larger $\beta$ yields a flatter, higher-entropy distribution, whereas smaller $\beta$ concentrates more sharply around the minimizers of $h$.
+Applying this operator to a possibly time-varying potential energy function $h(y\mid\mathcal{D}_{t})$ gives a preference distribution over outcomes, and hence
+$\begin{split}G&\approx-I(s;(x,y)\mid\mathcal{D}_{t})-\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[\log{e^{-h(y\mid\mathcal{D}_{t})/\beta}}]+Z\\ &=-I(s;(x,y)\mid\mathcal{D}_{t})+\frac{1}{\beta}\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[h(y\mid\mathcal{D}_{t})]+Z,\end{split}$
+where $q(y\mid x,\mathcal{D}_{t})$ is the predictive distribution of a surrogate model constructed from the historical data $\mathcal{D}_{t}$, and $Z=\log\int_{\mathcal{Y}}e^{-h(y\mid\mathcal{D}_{t})/\beta}dy$ is a normalization constant independent of $x$.
+This leads to our proposed decision principle:
+$x_{t+1}=\arg\max_{x\in\mathcal{X}}\{\beta_{t}\underbrace{I(s;(x,y)\mid\mathcal{D}_{t})}_{\text{epistemic}}-\underbrace{\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[h(y\mid\mathcal{D}_{t})]\}}_{\text{pragmatic}}.$
+Here, the conditional mutual information $I(s;(x,y)\mid\mathcal{D}_{t})$ captures the epistemic drive to reduce uncertainty about latent quantities of interest; the expected potential $\mathbb{E}[h(y\mid\mathcal{D}_{t})]$ captures the pragmatic drive to avoid undesirable outcomes; and $\beta_{t}\geq 0$ controls the trade-off between the two.
+By construction, this decision rule balances information-seeking and goal-directed behavior within a single objective.
+Rather than treating exploration and exploitation as competing heuristics, it expresses them as two inseparable aspects of one decision criterion: the agent seeks information insofar to improve outcome-sensitive action.
+In other words, it demonstrates a pragmatic curiosity (PraC).
+3.3 A Unified View of Acquisition Strategies in BO and BED
+PraC serves as a unifying umbrella principle for both learning-oriented and optimization-oriented acquisition strategies.
+By varying the epistemic target $s$ and the pragmatic potential $h_{t}(y)$, many classical acquisition rules from BO and BED can be recovered as special cases or close approximations, as summarized in Table 1.
 Table 1: A unified view of different acquisition strategies in BO and BED, where $x^{\ast}$, $y^{\ast}$ represent the true optimal solution and value, respectively, and $\hat{y}$ is the best value observed in $\mathcal{D}_{t}$.
-Acquisition Strategy Acquisition Function $\alpha(x)$ Expected Free Energy
-Parameters $s$ Preferences $p(y)$
-Expected Information Gain (Chaloner and Verdinelli, 1995) $I(\theta;(x,y)|\mathcal{D}_{t})$ $\theta$ -
-Entropy Search (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014) $I(x^{*};(x,y)|\mathcal{D}_{t})$ $x^{*}$ -
-Max-value Entropy Search (Wang and Jegelka, 2017) $I(y^{*};(x,y)|\mathcal{D}_{t})$ $y^{*}$ -
-Joint Entropy Search (Hvarfner et al., 2022b) $I((x^{*},y^{*});(x,y)|\mathcal{D}_{t})$ $(x^{*},y^{*})$ -
-Bayesian Algorithm Execution (Neiswanger et al., 2021) $I(\mathcal{O}_{\mathcal{A}}(f);(x,y)|\mathcal{D}_{t})$ $\mathcal{O}_{\mathcal{A}}(f)$ -
-GP-Upper Confidence Bound $\mu_{t}(x)+\beta^{1/2}\sigma_{t}(x)$ $f_{\mathcal{X}}$ $\exp\{y\}$
+Acquisition Strategy Acquisition Function Pragmatic Curiosity
+$s$ $h_{t}(y)$
+Expected Information Gain (Chaloner and Verdinelli, 1995b) $I(\theta;(x,y)\mid\mathcal{D}_{t})$ $\theta$ -
+Entropy Search (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014) $I(x^{*};(x,y)\mid\mathcal{D}_{t})$ $x^{*}$ -
+Max-value Entropy Search (Wang and Jegelka, 2017) $I(y^{*};(x,y)\mid\mathcal{D}_{t})$ $y^{*}$ -
+Joint Entropy Search (Hvarfner et al., 2022) $I((x^{*},y^{*});(x,y)\mid\mathcal{D}_{t})$ $(x^{*},y^{*})$ -
+Bayesian Algorithm Execution (Neiswanger et al., 2021) $I(\mathcal{O}_{\mathcal{A}}(f);(x,y)\mid\mathcal{D}_{t})$ $\mathcal{O}_{\mathcal{A}}(f)$ -
+GP-Upper Confidence Bound  [^1] Strictly speaking, the correspondence for GP-UCB is approximate rather than exact, because GP-UCB uses a first-order standard-deviation term rather than a second-order variance term.
+Detailed derivations are provided in Appendix B.
+Still, the connection reveals the information-theoretic structure underlying this heuristic form.
+(Srinivas et al., 2009) $\mu_{t}(x)+\beta^{1/2}\sigma_{t}(x)$ $f_{\mathcal{X}}$ y-y
 […]
 […]
-The reinterpretations of most of the acquisition strategies in Table 1 are straightforward according to their definitions.
-However, placing a rather intuitive GP-UCB strategy within this framework seems implicit.
-To reveal their connection, we rely on the following lemma:
-Lemma 3.2.
-Let $\mathbf{X}\subseteq\mathcal{X}$ be a subset of inputs, and the corresponding function values evaluated at those inputs be denoted as $f_{\mathbf{X}}$.
-Given a historical dataset $\mathcal{D}_{t}$, and new measurements $\mathbf{Y}$ observed at $\mathbf{X}$, the mutual information
-$I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D}_{t})=I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}_{t}),$
+More importantly, this unifying principle opens the door to acquisition design for a broader class of hybrid learning-and-optimization problems that go beyond classical BO and BED.
+Such problems are typically governed by three coupled meta-decisions: (i) what the agent seeks to know, (ii) what it seeks to achieve, and (iii) how it trades off those two.
+In the next section, we formalize these three design axes through symbols, regrets, and curiosity.
+4 The Triad of Meta-Decisions: Symbols, Regrets, and Curiosity
+One central contribution of PraC is that it foregrounds three meta-decisions underlying any hybrid learning-and-optimization problem.
+First, the agent must decide which latent quantity should be represented and clarified; this is implied by the choice of $s$.
+Second, it must decide how outcomes should be evaluated; this is encode through the choice of $h_{t}$.
+Third, it must decide how strongly unresolved uncertainty should influence action; this corresponds to the choice of $\beta_{t}$.
+These three choices are the critical operational levers through which an abstract decision paradigm becomes a practical decision rule.
+This section develops those three choices both conceptually and operationally.
+We formalize symbols to specify the latent quantities whose clarification can change what the agent ought to do, regrets to provide an operational loss landscape over outcomes, and curiosity to regulate the exchange rate between clarifying knowledge and pursuing goals.
+4.1 Symbols and Models
+“All models are wrong, but some are useful.” A model becomes useful when it represents the distinctions whose clarification can change what we should do. In PraC, we call these distinctions symbols.
+Different modeling regimes induce different forms of symbols.
+Parametric models.
+In parametric models, symbols are typically fixed global symbols: the latent state is represented by a finite-dimensional parameter or state variable whose components retain persistent semantic identities across decisions.
+Examples include unknown physical coefficients, latent modes, transition parameters, or class labels.
+In this regime, the epistemic objective is to reduce uncertainty about these stable hidden meanings.
+Non-parametric models.
+In non-parametric models, however, there may be no finite set of globally privileged coordinates.
+Instead, the relevant symbols are often induced local symbols, created by the contemplated observations themselves.
+This is formalized by the following representation-compression property.
+Lemma 4.1 (Representation Compression).
+Let $\mathbf{X}\subseteq\mathcal{X}$ be a subset of inputs, and let $f_{\mathbf{X}}$ denote the corresponding function values.
+Given a historical dataset $\mathcal{D}_{t}$ and new measurements $\mathbf{Y}$ observed at $\mathbf{X}$, then
+$I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})\mid\mathcal{D}_{t})=I(f_{\mathbf{X}};\mathbf{Y}\mid\mathcal{D}_{t}),$
 for any (finite or infinite) set $\mathcal{X}$.
 Proof.
-See Appendix C. ∎
-Then if we assume constant Gaussian noises $\mathcal{N}\sim(0,\sigma^{2})$ for the observations, we have
-$\begin{split}I(f_{x};y|\mathcal{D}_{t})&=H(y|\mathcal{D}_{t})-H(y|f_{x},\mathcal{D}_{t})\\ &=\frac{1}{2}\log(1+\sigma^{-2}\sigma_{t}^{2}(x)),\end{split}$
-where $\sigma_{t}^{2}(x)$ is the variance evaluated on the GP model $p(f_{x}|\mathcal{D}_{t})$.
-When further assuming that the GP kernel $\kappa_{t}(x,x^{\prime})\leq 1,\forall x,x^{\prime}\in\mathcal{X}$, then $0\leq\sigma_{t}^{2}(x)\leq\kappa_{t}(x,x^{\prime})\leq 1$, which gives
-$\log(1+\sigma^{-2}\sigma_{t}^{2}(x))\geq\log(1+\sigma^{-2})\sigma_{t}^{2}(x).$
-If we choose $\beta=\frac{1}{2}\log(1+\sigma^{-2})$, then the epistemic term in EFE, i.e., $I(f_{x};y|\mathcal{D}_{t})$, provides an upper bound of the square of the exploration term $\beta^{1/2}\sigma_{t}(x)$ in GP-UCB.
-This reveals the close relationship between GP-UCB and AIF, showing that even a seemingly pure intuitive strategy, like GP-UCB, can have rather rigorous mathematical foundations underlying it.
-4 A New Paradigm to Derive Acquisition Functions
-The preference distribution $p(y)$ in Table 1 can be interpreted as a softmax transformation of a time-varying value function over the outcomes.
-Generalizing from this observation, we propose a new paradigm for deriving acquisition functions applicable to a wider class of hybrid problems other than traditional BO and BED.
-Grounded on Boltzmann distribution (also called Gibbs distribution) in statistical mechanics, we define an Boltzmann operator $\mathcal{B}_{\beta}$ that maps an energy function $h(z)$ over domain $\mathcal{Z}$ into a Boltzmann distribution:
-$(\mathcal{B}_{\beta}[h])(z)\coloneq\frac{e^{-h(z)/\beta}}{\int_{\mathcal{Z}}e^{-h(z)/\beta}dz},$
-where $\beta$ is called a temperature parameter (in allusion to statistical mechanics).
-A higher temperature results in a more uniform output distribution (i.e.“more random” with higher entropy), while a lower temperature results in a sharper output distribution, with the maximizers of $h(z)$ dominating.
-Then, for any pre-defined energy function $h(y|\mathcal{D}_{t})$, we can transfer it into a Boltzmann distribution and use it as the preference distribution $p(y)$ in (3):
-$\begin{split}G&=-I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\log{e^{-h(y|\mathcal{D}_{t})/\beta}}]+Z\\ &=-I(s;(x,y)|\mathcal{D}_{t})+\frac{1}{\beta}\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]+Z,\end{split}$
-where $p(y|x,\mathcal{D}_{t})$ is the predictive distribution of a surrogate model constructed from the historical data $\mathcal{D}_{t}$, and $Z=\log\int_{\mathcal{Y}}e^{-h(y|\mathcal{D}_{t})/\beta}dy$ is a normalization constant independent of $y$.
-Therefore, we propose a new acquisition paradigm, pragmatic curiosity, by minimizing the EFE, i.e.,
-$x_{t+1}=\arg\max_{x\in\mathcal{X}}\{\beta_{t}I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]\},$
-where $\beta_{t}\geq 0$ is a parameter that modulates the degree of curiosity, and $h(y|\mathcal{D}_{t})$ is a problem-dependent energy function that captures the goal (in terms of regret w.r.t.
-a certain outcome $y$).
-The curiosity $\beta_{t}$ plays an important role in balancing the performance between learning and optimization.
-Li et al.
-(2026) derive a formal theoretical guarantee of posterior consistency and bounded cumulative regret under a sufficient curiosity condition, and provide practical design guidelines for choosing $\beta_{t}$ in different settings.
-This paper instead examines how a problem-dependent energy function, $h(y|\mathcal{D}_{t})$, provides a flexible mechanism for explicitly representing uncertainty over the goals themselves.
-This added expressiveness enables acquisition design for a broad class of complex settings that standard methods often overlook, including tasks with evolving goals (where conditions change over time) and implicit goals (not specified a priori).
-The effectiveness and versatility of this formulation on such tasks are demonstrated in the next section.
+See Appendix A. ∎
+Lemma 4.1 shows that, under a non-parametric model such as a Gaussian process, the information gained from observing $\mathbf{Y}$ at $\mathbf{X}$ is mediated entirely through the function values $f_{\mathbf{X}}$ that directly generate those observations.
+These local evaluations play a role analogous to parameters in parametric models or latent states in partially observed dynamical systems, but they need not belong to a fixed global representation.
+This compression reveals both the power and the cost of non-parametric representation.
+The power is parsimony: the agent need not reason about every component of an infinite-dimensional latent object when only a local slice can influence the prospective observation.
+The cost is semantic instability: because the symbol is query-dependent, its meaning is induced by the decision context rather than fixed in advance.
+This also suggests a broader role for PraC.
+PraC asks not merely how expressive a model is, but whether the symbol used by the acquisition is sufficient for action evaluation, minimal enough to avoid epistemic redundancy, and structured enough to support interpretable decision-making.
+Thus beyond selecting actions, PraC can also guide the construction of representations.
+When the relevant structure is unclear, induced local symbols can reveal which distinctions are repeatedly useful for decision.
+Over time, such recurring local distinctions may be consolidated into more stable global symbols.
+In this sense, PraC does not only act within a representation; it also provides a criterion for which representations deserve to persist.
+4.2 Regrets and Goals
+“All successes are successful alike, all failures are failed in their own way.” Most of the time, we don’t know what the goal is until we arrive there.
+It seems easier to address what it is not than what it is: success is often rare and difficult to characterize in full, whereas shortfall, inconsistency, or violation is much more common and easier to recognize and certify.
+For this reason, it is often more natural to encode what an outcome fails to achieve than to assume that a complete reward landscape is already known.
+Under this view, we define regret as the shape cast by a goal onto outcome space.
+Regret thus acts as a carrier of goal: it gives an operational representation of what matters by quantifying how an outcome falls short of what is desired.
+We begin by formally defining a generalized notion of regret.
+Definition 4.2 (Regret Function).
+Let $\mathcal{Y}$ denote the outcome space, and let $\xi$ denote a desired reference, such as an optimal outcome, a target set, a safety specification, or a preference criterion.
+A regret function is a mapping
+$r(\cdot;\xi):\mathcal{Y}\to\mathbb{R}_{\geq 0}$
+that quantifies how an outcome $y$ falls short of the desired reference $\xi$.
+It obtains 0 for the desired outcomes, i.e., $\inf_{y\in\mathcal{Y}}r(y;\xi)=0$, with smaller values indicating outcomes that are closer to what is desired.
+This abstraction recovers several important cases:
+  • When $\xi=y^{\ast}$ is the global optimum of an objective function, $r(y;\xi)$ reduces to the conventional instantaneous regret in BO.
+  • When no outcome-dependent pragmatic preference is imposed, as in pure BED, one may take $r(y;\xi)\equiv 0$, so that decision-making is driven entirely by epistemic value.
+  • Intermediate choices recover hybrid objectives that jointly encode performance, safety, feasibility, or other task-specific desiderata.
+In practice, however, the ideal regret function $r(\cdot;\xi)$ is generally unknown, implicit, or computationally intractable.
+A decision-maker therefore cannot act directly on $r$.
+Instead, it must rely on a belief-conditioned surrogate that reflects its current internal assessment of outcome desirability.
+This motivates the following definition of a potential energy function.
+Definition 4.3 (Potential Energy Function).
+The potential energy function is a belief-conditioned surrogate regret
+$h_{t}:\mathcal{Y}\to\mathbb{R}_{\geq 0},\qquad h_{t}(y):=h(y\mid\mathcal{D}_{t}),$
+that represents the agent’s current internal assessment of the undesirability of outcome $y$ based on the information available up to time $t$.
+Without loss of generality, we may shift $h_{t}$ so that $\inf_{y\in\mathcal{Y}}h_{t}(y)=0.$ Thus, $h_{t}$ serves as a time-varying, belief-dependent approximation of the true regret structure, and is updated recursively as new data are acquired.
+The distinction between $r$ and $h_{t}$ is what makes PraC operational.
+The regret function $r$ represents the ideal task-level notion of how an outcome falls short of what is desired, whereas $h_{t}$ is the best currently available decision-usable surrogate induced by the agent’s present belief state.
+This mirrors the epistemic approximation used earlier in deriving PraC: just as the inaccessible true posterior is replaced by the current surrogate belief, the inaccessible true regret landscape is replaced by a belief-conditioned potential over outcomes.
+As more data are collected, both the predictive model and the potential energy evolve, so that the agent’s internal representation of desire is refined together with its understanding of the world.
+This distinction also clarifies how $h_{t}$ should be designed in different regimes.
+Known goals.
+When the goal is known, $h_{t}$ can be designed to directly encode domain knowledge.
+In this case, the designer already has access to a meaningful notion of what constitutes shortfall, so the potential energy can be specified explicitly.
+Examples include classical regret with respect to an optimum, distance to a target set, penalties for constraint violation, or weighted combinations of task performance and safety requirements.
+In this regime, the role of $h_{t}$ is to express a known desiderative structure in a form that can be evaluated under the predictive model.
+Unknown goals.
+When the goal is unknown or only partially specified, the regret itself must be learned.
+In this regime, the pragmatic term should be endowed with additional structure, so that the agent can infer not only how the world behaves, but also what outcomes should count as desirable or undesirable.
+A natural construction is hierarchical: outcome desirability is mediated by additional latent variables or higher-level preference models.
+Given an unknown regret model $g:\mathcal{Y}\to\mathbb{R}_{\geq 0}$, one can define $h_{t}$ through an inner EFE over that regret model, yielding the nested acquisition
+[…]
+where $\beta_{t},\gamma_{t}\geq 0$, and $z\sim g(y)$ encodes a second-order regret, namely the regret assigned by the learned regret model itself.
+In this way, the agent does not optimize with respect to a fully specified reward function given in advance; rather, it learns a surrogate of regret jointly with the predictive model.
+This yields a richer and more flexible account of the interplay between learning and optimization, and enables multiple layers of inference to be composed within a single decision rule.
+For instance, as we show later in 5.3, the model $g(y)$ can be learned through preference feedback.
+This point highlights a key advantage of the PraC paradigm.
+Unlike many mainstream approaches that presuppose a fully specified reward function, PraC requires only a current surrogate of what can already be recognized as desirable or regrettable, together with a mechanism for refining that surrogate as evidence accumulates.
+One need not begin with complete illumination; it is enough to start by seeing a light in the darkness, and to let both understanding of the world and understanding of value sharpen together.
+4.3 The Degree of Curiosity
+If symbols specify what hidden distinctions are worth learning, and regrets specify what desired outcomes are worth pursuing, then the coefficient $\beta_{t}$ determines how much immediate pragmatic sacrifice the agent is willing to tolerate in order to refine its epistemic understanding.
+In this sense, $\beta_{t}$ is the exchange rate between clarifying knowledge and pursuing goals, which we denote as the degree of curiosity.
+Different regimes of $\beta_{t}$ produce different modes of behavior.
+When $\beta_{t}=0$, the agent treats its current symbols as fixed and acts myopically with respect to the present surrogate of regret.
+At the other extreme, a very large $\beta_{t}$ prioritizes information acquisition even when that information has weak pragmatic relevance.
+Neither extreme is satisfactory: seeking knowledge without goals degenerates into aimless information gathering, whereas pursuing goals without clarified meanings risks acting on an incomplete understanding of the world.
+This reveals a closed-loop structure between knowledge and action.
+The agent acts according to its current symbolic understanding and regret surrogate; the resulting observation then reshapes that understanding and changes which future actions become desirable.
+Thus, curiosity should not be viewed as a static exploration heuristic.
+Rather, it is a regulated gain in this knowledge–action loop: it determines how strongly unresolved uncertainty should influence action when the agent’s current understanding may still be insufficient for reliable decision-making.
+This motivates designing $\beta_{t}$ as an adaptive controller for the knowledge–action loop.
+Feedforward scale.
+The feedforward component estimates the scale of curiosity required for epistemic value to be comparable with pragmatic value. Given a finite candidate set $\mathcal{X}_{t}^{\mathrm{cand}}$, let
+$\widehat{I}_{t}(x)\approx I(s;(x,y)\mid\mathcal{D}_{t}),\qquad\widehat{H}_{t}(x):=\mathbb{E}[h_{t}(y)\mid x,\mathcal{D}_{t}],$
+where $\widehat{I}_{t}(x)$ estimates the epistemic value of querying $x$, and $\widehat{H}_{t}(x)$ estimates its expected pragmatic regret.
+To avoid unstable ratios when information gain is negligible, define the informative candidate set
+$\mathcal{X}_{t}^{\mathrm{info}}=\{x\in\mathcal{X}_{t}^{\mathrm{cand}}:\widehat{I}_{t}(x)>\varepsilon_{I}\},$
+where $\varepsilon_{I}>0$ is a numerical floor. If $\mathcal{X}_{t}^{\mathrm{info}}$ is nonempty, we define the feedforward curiosity scale as
+$\beta_{t}^{\mathrm{ff}}=\operatorname{Quantile}_{x\in\mathcal{X}_{t}^{\mathrm{info}}}\frac{\widehat{H}_{t}(x)}{\widehat{I}_{t}(x)+\varepsilon_{I}}.$
+If $\mathcal{X}_{t}^{\mathrm{info}}=\varnothing$, we set $\beta_{t}^{\mathrm{ff}}=\beta_{\min}$, since the current candidate set contains no sufficiently informative query.
+This feedforward term $\beta_{t}^{\mathrm{ff}}$ estimates the local exchange rate between information and regret: when useful information is scarce relative to pragmatic regret, the required curiosity scale increases; when information is readily available, a smaller scale is sufficient to make information compete with immediate pragmatic value.
+Feedback activation.
+The feedback component determines whether epistemic pressure is still needed.
+Let $U_{t}$ be a scalar uncertainty measure over the relevant symbol, such as posterior entropy, and let $U^{\star}$ denote the desired epistemic equilibrium, typically zero or an irreducible uncertainty floor.
+Define
+$\beta_{t}^{\mathrm{fb}}=k_{\beta}\operatorname{clip}\left(\frac{U_{t}-U^{\star}}{U_{0}-U^{\star}+\varepsilon_{U}},0,1\right),$
+where $U_{0}$ is the initial uncertainty, $k_{\beta}$ is a feedback gain, and $\varepsilon_{U}>0$ prevents numerical instability.
+The combined schedule for curiosity coefficient is then
+$\beta_{t}=\operatorname{clip}\left(\beta_{t}^{\mathrm{fb}}\beta_{t}^{\mathrm{ff}},\beta_{\min},\beta_{\max}\right),$
+where the two components play complementary roles: the feedforward term estimates how large curiosity must be for information to compete with pragmatic value, while the feedback activation suppresses curiosity as the belief approaches the desired epistemic equilibrium.
+Thus, curiosity is high only when information is both decision-relevant and epistemically needed; it relaxes when the relevant uncertainty has been resolved.
 5 Experiments
-In this section, we exemplify the benefits and variability of the proposed acquisition strategy by performing experiments across three categories that differ substantially from one another.
-These problems originate from distinct literature within BO and BED, and none of them are canonical BO or BED tasks.
-Thus each of them is evaluated against a different set of BO-type (optimization-focused) and BED-type (learning-focused) baselines appropriate for that specific task, ensuring a fair and rigorous comparison.
-Figure 1: Performance evaluation for constrained system identification on environmental monitoring in 2d plume fields. Error bars represent $\pm 1$ std over 5 seeds.
-5.1 Parametric Models with Known Invariant Goals
-We first consider a simplest task where the goals are known and invariant, and we have a decent understanding of the model, where only certain finite parameters are unknown.
-One typical problem structure is the constrained system identification whose objective is to precisely learn unknown system parameters, $\theta$, when valid observations can only be gathered under specific operational constraints, quantified by $C(y)\leq 0$.
-For instance, when monitoring a chemical plume, the goal is to identify its source ($\theta$) while ensuring that sensor measurements ($y$) do not exceed a saturation threshold ($C(y)$).
-In such cases, it is straightforward to choose the interested state as $s=\theta$, and the energy function as $h(y|\mathcal{D}_{t})=\mathbb{I}(C(y)>0)$. This leads to an acquisition function:
-$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(\theta;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}\left[\mathbb{I}(C(y)>0)\right].\end{split}$
+PraC provides a principled but flexible acquisition-design template through three coupled choices: a symbolic representation of what is worth learning, a regret-based potential that encodes what outcomes matter, and a curiosity coefficient that determines how strongly epistemic and pragmatic pressures are traded against one another.
+In this section, we instantiate these choices across three regimes of increasing complexity.
+The experiments are organized according to the structure exposed by PraC.
+We begin with fixed global symbols and known goals, where the latent quantity is a finite hypothesis and the regret is a downstream Bayes risk.
+We then consider induced local symbols with known but evolving goals, where the relevant latent quantity is induced by a non-parametric surrogate and the goal depends on the coverage already obtained.
+Finally, we study a hierarchical setting with unknown goals, where the agent must learn both how actions produce outcomes and how those outcomes should be valued.
+We study dynamic scheduling of $\beta_{t}$ in detail in the first regime, where its closed-loop interpretation is most transparent.
+In the non-parametric and hierarchical regimes, we use fixed curiosity weights to isolate the effects of representation and regret design.
+Figure 1: Decision-oriented plume-monitoring performance over 20 seeds.
+Top row: posterior Bayes risk.
+Bottom row: task-specific realized decision metric: response loss for source response localization and consequence-weighted dispatch, and missed-source risk for active source prioritization.
+Shaded regions show mean $\pm$ standard error.
+5.1 Fixed Global Symbols with Known Goals
+We begin with the most transparent regime, in which the symbolic structure is fixed and the downstream goal is explicitly specified.
+Here the unknown quantities are finite-dimensional hypotheses with stable semantic meaning across all decisions, so the epistemic target is a fixed global symbol.
+At the same time, the task objective is known in advance and can be expressed through a closed-form downstream regret.
+This regime provides the cleanest setting in which to examine whether curiosity can be elevated from a heuristic coefficient to a monitored and interpretable part of the decision process.
+A representative problem of this form is decision-oriented monitoring, where the purpose of sensing is not merely to estimate environmental parameters, but to improve a downstream response decision.
+At each iteration, the agent maintains a posterior over latent environmental hypotheses $\theta$, selects a sensor query $x$, observes a plume count $y$, updates the posterior, and evaluates the downstream decision induced by the updated belief.
+Let $a\in\mathcal{A}$ denote a downstream action, and let $L(a,\theta)$ denote the task-specific loss incurred by taking action $a$ when the true hypothesis is $\theta$.
+Given the current surrogate posterior $q_{t}(\cdot)$, the Bayes risk is
+$\mathrm{BR}(\mathcal{D}_{t})\coloneq\min_{a\in\mathcal{A}}\mathbb{E}_{\theta\sim q_{t}}\left[L(a,\theta)\right].$
+This measures the downstream loss the agent still expects to incur if it acts optimally under its current belief.
+For a candidate sensor query $x$, define the expected Bayes-risk reduction
+$\widehat{\Delta\mathrm{BR}}_{t}(x)\coloneq\mathrm{BR}(\mathcal{D}_{t})-\mathbb{E}_{q_{t}(y\mid x)}\left[\mathrm{BR}(\mathcal{D}_{t}\cup\{(x,y)\})\right].$
+The PraC acquisition becomes
+$\alpha_{t}(x)=\beta_{t}I(\theta;(x,y)\mid\mathcal{D}_{t})+\widehat{\Delta\mathrm{BR}}_{t}(x).$ (4)
 Tasks.
-We perform experiments on a real-world environmental monitoring problem in 2d plume fields where the sensors have a saturation threshold $y_{max}$ (i.e., $C(y)=y-y_{max}$) (Detailed settings and hyper-parameter choices in Appendix D.2).
-We consider three types of monitoring tasks: (a) locating the unknown source location; (b) estimating unknown wind direction and strength; and (c) identifying the active sources in fields with multiple sources.
+We evaluate this setting on three decision-oriented plume-monitoring tasks in 2D fields.
+(a) Source response localization asks the agent to localize a plume source well enough to dispatch a response team near it.
+Here the downstream action $a$ is a response location, and the loss $L(a,\theta)$ is the normalized squared distance between the response location and the source location specified by $\theta$.
+(b) Consequence-weighted dispatch uses the same response-decision structure, but errors in high-consequence regions are assigned larger loss.
+(c) Active source prioritization asks the agent to select which suspected industrial leaks to repair under a limited crew budget.
+Here the downstream action $a$ is a subset of sources, and the loss $L(a,\theta)$ is the weighted risk of true active sources not selected for repair.
+Detailed plume models, hypothesis spaces, downstream losses, and Bayes-risk computations are provided in Appendix C.2.
 Baselines.
-We compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by choosing the point that leads to the least probability of violating the constraint (BO-type); and (c) EIG about the unknown parameter (BED-type).
-Evaluations. We evaluate the performance from both epistemic and pragmatic perspectives: (a) estimation accuracy; and (b) constraint violations.
+We compare two variants of PraC with (i) dynamic $\beta$ as scheduled in Section 4.3, and (ii) fixed-$\beta$ with $\beta=5$, against (a) random sampling, (b) pure information gain (EIG), and (c) decision-greedy policy by minimizing $BR_{t}(x)$.
+Evaluations.
+We report posterior Bayes risk and task-specific realized decision metrics.
+For source response localization and consequence-weighted dispatch, the realized metric is response loss under the true source.
+For active source prioritization, the realized metric is missed-source risk, namely the total consequence weight of true active sources not selected for repair.
+Additional metrics, including response success, parameter error, top-$k$ recall, weighted top-$k$ recall, and detailed analysis are reported in Appendix C.2.6.
 Results.
 Fig.
-1 shows that our approach achieves consistently stronger query efficiency over baselines while respecting all operational constraints, with cumulative constraint violation always being zero.
-This advantage is especially evident in the source localization task, where the drive for information and the need to satisfy constraints create opposing pressures.
-Resolving this conflict, our approach reaches near-perfect estimation using up to 40% fewer queries than competing methods.
-Figure 2: Performance evaluation for targeted active search on failure discovery in autonomous driving scenarios. Error bars represent $\pm 1$ std over 4 seeds.
-5.2 Non-Parametric Models with Known Evolving Goals
-Next, we consider a more challenging setting where the task condition evolves, and the model is fully black-box such that we need to resort to a non-parametric model (e.g., GPs).
-One such example is the targeted active search in multi-objective design problems.
-The objectives are viewed as metrics where specific ranges carry particular significance, and the goal is to design experiments that maximize coverage of these important regions $\mathcal{S}$.
-Adopted from Malkomes et al.
-(2021), we assume that the experimental design problems possess a sense of known resolution, such as simulation accuracy or manufacturing precision/tolerance, and any outcome within distance $\delta$ of another does not convey extra information about $\mathcal{S}$.
-Similar to Malkomes et al.
-(2021), we define the coverage neighborhood of any $y$ as $\mathbb{C}_{\delta}(y):=\{y^{\prime}:d(y,y^{\prime})<\delta\}$.
-And the coverage neighborhood of a set of points $Y$ is defined as $\mathbb{C}_{\delta}(Y):=\bigcup_{y\in Y}\mathbb{C}_{\delta}(y)$.
-Since the goal is to cover as much volume of the interested region $\mathcal{S}$ as possible, $h(y|\mathcal{D}_{t})$ can be chosen as $h(y|\mathcal{D}_{t})=\text{Vol}(\mathcal{S})-\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})$, where $Y$ contains all the history of outcomes.
-Similarly, coverage over the design configurations can be flexibly incorporated by expanding the interested state into the whole input space, i.e., $s=f_{\mathcal{X}}$.
-With non-parametric models like GPs, we can leverage Lemma 3.2 and derive an acquisition function:
-$\begin{split}\alpha(x|\mathcal{D}_{t})=\beta I(f_{x};y|\mathcal{D}_{t})+\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\text{Vol}(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S})].\end{split}$
+1 shows that dynamic PraC gives the strongest overall performance across the three decision-oriented plume tasks.
+In source response localization, dynamic PraC achieves the lowest final Bayes risk while matching the best realized response loss.
+In consequence-weighted dispatch, decision-greedy obtains the lowest final Bayes risk, but dynamic PraC obtains the best realized response loss and the lowest parameter error.
+This distinction is important: Bayes risk measures expected loss under the agent’s current posterior, whereas realized response loss evaluates the downstream decision under the true latent environment.
+Dynamic curiosity can therefore sacrifice some immediate posterior-risk reduction in order to collect information that improves the realized downstream decision.
+In active source prioritization, dynamic PraC obtains the lowest final Bayes risk, improving over fixed-$\beta$ PraC, decision-greedy, EIG, and random sampling.
+Dynamic and fixed-$\beta$ PraC tie on missed-source risk, while both substantially improve over EIG and random sampling.
+Dynamic scheduling of curiosity.
+Scheduler diagnostics are reported in Appendix C.2.7.
+The feedforward scale estimates the local exchange rate between Bayes-risk reduction and information gain, while the feedback activation suppresses curiosity as decision-relevant uncertainty decreases.
+These diagnostics support the intended interpretation of $\beta_{t}$: dynamic PraC does not merely reduce environmental uncertainty for its own sake, but learns distinctions that matter for downstream response decisions.
+Figure 2: Performance evaluation for targeted active search on failure discovery in autonomous driving scenarios.
+Top row: coverage in the outcome/metric space.
+Bottom row: coverage in the input/parameter space.
+Error bars represent $\pm 1$ standard deviation over 4 seeds.
+5.2 Induced Local Symbols with Known Evolving Goals
+We now move to a more challenging regime in which the latent structure is no longer naturally captured by a fixed finite-dimensional parameter.
+Instead, the relevant symbolic representation is induced locally by the contemplated observations, as described in Section 4.1.
+At the same time, the goal remains known, but it is no longer invariant: it evolves with the outcomes already collected.
+This regime is designed to highlight the value of PraC’s symbolic flexibility.
+To isolate that effect, we retain a fixed curiosity weight in this subsection.
+A representative problem of this form is targeted active search in multi-objective design.
+Here the objective is not simply to optimize a scalar reward, but to cover as much as possible of an outcome region $\mathcal{S}$ deemed important.
+Following Malkomes et al.
+(2021b), we assume a known resolution $\delta$, so that outcomes within distance $\delta$ are treated as redundant with respect to target coverage.
+Define the coverage neighborhood of an outcome $y$ and a set of outcomes $Y$ by
+$\mathbb{C}_{\delta}(y):=\{y^{\prime}:d(y,y^{\prime})<\delta\},\qquad\mathbb{C}_{\delta}(Y):=\bigcup_{y\in Y}\mathbb{C}_{\delta}(y).$
+Then a natural potential is
+$h_{t}(y)=\mathrm{Vol}(\mathcal{S})-\mathrm{Vol}\!\left(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S}\right),$
+which penalizes outcomes that add little new coverage of the target set.
+Meanwhile, the relevant epistemic target is the local non-parametric symbol induced by the contemplated observation, so we take $s=f_{\mathcal{X}}$ and use the representation-compression result in Lemma 4.1 to obtain
+$\alpha(x\mid\mathcal{D}_{t})=\beta I(f_{x};y\mid\mathcal{D}_{t})+\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}\!\left[\mathrm{Vol}\!\left(\mathbb{C}_{\delta}(Y\cup y)\cap\mathcal{S}\right)\right].$
 Tasks.
-We perform experiments on a real-world failure discovery problem in autonomous driving scenarios where the perception module, a YOLO detector, may fail due to multiple causes, which could potentially lead to collisions (3d input-2d output).
-We consider three target sets with decreasing volume, i.e., Target Set 1 $\supset$ Target Set 2 $\supset$ Target Set 3 (Detailed settings and hyper-parameter choices in Appendix D.3).
+We study failure discovery for a YOLO-based perception module in autonomous-driving scenarios.
+The input is a 3-dimensional scenario parameterization, and the outcome is a 2-dimensional failure metric describing two failure modes that may lead to collision.
+We consider three nested target sets of decreasing volume, $\mathcal{S}_{1}\supset\mathcal{S}_{2}\supset\mathcal{S}_{3},$ where smaller target sets correspond to rarer and more difficult failure regions.
+Details of the CARLA simulation, failure metrics, and target-set definitions are provided in Appendix C.3.
 Baselines.
-We again compare our proposed acquisition strategy (AIF) with BO-type and BED-type baselines tailored this task: (a) Random; (b) Greedy by maximizing coverage volume in metrics space (BO-type); and (c) EIG by maximizing coverage volume in parameter space (BED-type).
-Evaluations. We evaluate the performance by inspecting the coverage volume over both spaces: (a) metrics space ($\mathbb{C}_{\delta}(Y)$); and (b) parameter space ($\mathbb{C}_{\delta}(X)$).
+We compare PraC against random sampling and two expected coverage improvement (ECI) variants adapted from Malkomes et al.
+(2021b): (a) ECI-y, which greedily prioritizes target coverage in the outcome/metric space, and (b) ECI-x, which prioritizes coverage in the input/parameter space.
+These baselines isolate the two coverage pressures that PraC jointly coordinates.
+Evaluations. We evaluate coverage in both spaces, including metric-space coverage $\mathbb{C}_{\delta}(Y)$, and parameter-space coverage $\mathbb{C}_{\delta}(X)$.
 Results.
 As shown in Fig.
-2, our AIF algorithm effectively balances coverage of both the parameter and metrics spaces.
-This capability is especially impactful for smaller target sets where the search is more difficult.
-In the most challenging case (Target Set 3), our method identifies nearly 10% more of the critical failure region compared to the leading baseline.
-Figure 3: Performance evaluation for composite BO with unknown preferences. Error bars represent $\pm 1$ std over 20 seeds for vehicle safety, penicillin, and 5 for energy resource allocation.
-5.3 Non-Parametric Models with Unknown Goals
-Finally, we look into the most difficult setting where both the models and goals are black-box.
-One practical scenario arises from the composite Bayesian optimization in multi-objective optimization problems.
-The objectives are weighted by a preference function $g(y)$ that is unknown a priori and must be simultaneously learned during the optimization process.
-As assumed in Lin et al.
-(2022), we can estimate the unknown preference function $g(y)$ by asking a decision-maker (DM) to express preferences over pairs of outcomes $(y_{1},y_{2})$.
-Let $z(y_{1},y_{2})\in\{1,2\}$ indicate whether the DM preferred the first or second outcome offered.
-Following Chu and Ghahramani (2005), we assume that the DM’s responses are distributed according to a probit likelihood: $L(z(y_{1},y_{2})=1|g(y_{1}),g(y_{2}))=\Phi(\frac{g(y_{1})-g(y_{2})}{\sqrt{2}\lambda}),$ where $\lambda$ is a hyperparameter, and $\Phi$ is the standard normal CDF.
-Taking the joint exploration and exploitation of the preference function $g(y)$ into consideration, we can choose $h(y|\mathcal{D}_{t})$ as the EFE of the preference model, i.e., $h(y|\mathcal{D}_{t})=\text{EFE}(g(y))$, which naturally extends the acquisition function into a nested structure:
-[…]
-where $x=[x_{1},x_{2}]$ are two jointly evaluated candidates, and $\gamma,\beta\geq 0$.
+2, PraC effectively balances the two forms of coverage.
+It does not merely spread samples broadly in parameter space, nor does it myopically chase already discovered target regions in outcome space.
+Instead, the information term encourages learning of the local surrogate structure, while the coverage potential directs samples toward still-uncovered parts of the target region.
+This coordination is especially useful for smaller target sets, where informative failures are sparse and purely greedy coverage strategies become brittle.
+In the most challenging case, Target Set 3, PraC discovers nearly 10 percentage points more normalized target-region coverage than the strongest baseline.
+Role of induced symbols.
+This experiment emphasizes the representational role of PraC.
+Once the epistemic target is allowed to be induced locally by the contemplated observation, the same acquisition principle naturally adapts from finite-hypothesis parameter identification to black-box target search with non-parametric surrogates.
+Figure 3: Performance evaluation for composite BO with unknown preferences.
+Metric is the best preference value attained among all collected outcomes.
+Error bars represent $\pm 1$ standard deviation over 20 seeds.
+Figure 4: Comparison with BOPE-style stage-wise variants for energy resource allocation. Metric is the best preference value attained. Error bars represent $\pm 1$ standard deviation over 20 seeds.
+5.3 Hierarchical Symbols with Unknown Goals
+Finally, we consider the most difficult regime, in which neither the predictive structure nor the goal is fully specified in advance.
+In this setting, the agent must learn not only how actions produce outcomes, but also how outcomes should be valued.
+This experiment is designed to emphasize the role of regret in PraC: the pragmatic term is no longer specified directly, but is itself inferred through an additional layer of modeling.
+To keep the focus on this hierarchical regret design, we retain fixed curiosity weights in this subsection.
+A representative example is composite Bayesian optimization with unknown preferences.
+The outcomes are multi-objective vectors, but the scalar preference over them is determined by an unknown function $g(y)$ that must be learned during optimization.
+Following Lin et al.
+(2022), we infer $g(y)$ from pairwise preference queries.
+For a pair of outcomes $\tilde{y}=[y_{1},y_{2}]$, let $l(y_{1},y_{2})\in\{1,2\}$ denote which outcome is preferred by the decision-maker.
+Following Chu and Ghahramani (2005), we assume the probit likelihood
+$\Pr(l(y_{1},y_{2})=1\mid g(y_{1}),g(y_{2}))=\Phi\!\left(\frac{g(y_{2})-g(y_{1})}{\sqrt{2}\lambda}\right),$
+where $\lambda$ is a hyperparameter and $\Phi$ is the standard normal CDF. Under this formation, outcome $y_{1}$ is more likely to be preferred when it has lower latent regret than $y_{2}$.
+At each iteration, the method selects a candidate pair $\tilde{x}=[x_{1},x_{2}]$ and observes the corresponding pair of outcomes $\tilde{y}=[y_{1},y_{2}]$ and a preference label $l$.
+This yields the hierarchical PraC acquisition
+$\begin{split}\alpha(\tilde{x}\mid\mathcal{D}_{t})=\beta_{t}I(f_{\mathcal{X}};(\tilde{x},\tilde{y})\mid\mathcal{D}_{t})+\mathbb{E}_{q(\tilde{y}\mid\tilde{x},\mathcal{D}_{t})}\left[\gamma_{t}I(g_{\mathcal{Y}};(\tilde{y},l)\mid\mathcal{D}_{t})-\min_{y\in\{y_{1},y_{2}\}}\mathbb{E}_{q(g_{y}\mid y,\mathcal{D}_{t})}[g_{y}]\right],\end{split}$ (5)
+where the first term learns the outcome model $f$; the second term learns the regret model $g$ through informative pairwise comparisons; and the final term exploits the currently predicted best member of the evaluated pair by favoring low expected regret.
 Tasks.
-We perform experiments on three real-world problems, including vehicle safety (5d-3d), penicillin production simulator (7d-3d), and distributed energy resource allocation in power grids (40d-4d) (Detailed settings and hyper-parameter choices in Appendix D.4).
+We evaluate this setting on three real-world multi-objective optimization problems of increasing complexity: (a) vehicle safety with 5-dimensional inputs and 3-dimensional outcomes, (b) penicillin production with 7-dimensional inputs and 3-dimensional outcomes, and (c) energy allocation in power grids with 40-dimensional inputs and 4-dimensional outcomes.
+Details of the simulators, outcome definitions, and ground-truth preference models are provided in Appendix C.4.
 Baselines.
-Different from BOPE in Lin et al.
-(2022), which separates and iterates the preference exploration and experimentation stages, our approach unifies the exploration and exploitation of both outcome and preference models simultaneously.
-In its best setting (i.e., switch stages and update models at each step), their acquisition strategy reduces to a special case of our proposed (4) with $\beta=\gamma=0$.
-Thus, we design a set of baselines to investigate the effect that each component in (4) plays for the joint exploration and exploitation of hierarchical models.
-We ablate the full AIF strategy (g + g IG + f IG) by (a) removing the last term (EIG); (b) removing the first term (g + g IG); and (c) removing the first two terms (g only).
-Evaluations. We evaluate their performance by inspecting the best preference of all collected outcomes using the true preference function $g(y)$.
+We first compare the full hierarchical PraC acquisition against ablated variants obtained by removing: (a) the outer epistemic term over $f$, (b) the inner epistemic term over $g$, or (c) both epistemic terms, leaving only the pragmatic preference term.
+These ablations isolate the contribution of outcome-model learning, preference-model learning, and their joint interaction.
+We additionally compare against BOPE-style stage-wise variants (Lin et al., 2022) for the energy-allocation task, where the distinction between joint and staged learning is most pronounced.
+Detailed design of BOPE-style variants are provided in Appendix C.4.3.
+Evaluations. We evaluate the best preference value attained among all collected outcomes using the true latent preference function.
 Results.
 Fig.
-3 illustrates our AIF approach’s outstanding ability to learn unknown preference functions, a key advantage over baselines that often fail due to ill-directed queries.
-As the tasks become more complex and noisy (from (a) to (c)), our results demonstrate that every component of our acquisition function (4) plays an irreplaceable role in achieving optimal performance.
-This benefit is most stark in the energy resource allocation task, where competing methods failed to capture any meaningful preference model while our approach consistently succeeded.
+3 shows that PraC consistently outperforms the ablated variants in learning and exploiting the unknown preference structure.
+As the tasks become more complex and noisier, each component of the hierarchical acquisition becomes increasingly important.
+This is most evident in the energy-allocation problem, where variants that remove either outcome-model exploration or preference-model exploration fail to discover high-preference regions reliably, whereas the full PraC acquisition continues to improve.
 Benefits of joint learning and optimization.
-To highlight the benefits of jointly learning and optimizing, rather than separating these into stages, we compare our method against several BOPE variants from Lin et al.
-(2022) that use different stage-wise design choices.
-Detailed experimental setups, design choices, plots, and analyses are provided in Appendix D.4.3.
-The results show that our method naturally balances exploration and exploitation at each step and consistently discovers higher-preference regions, whereas the BOPE variants are highly sensitive to how the stages are configured.
-Consequently, stage-wise approaches like BOPE require careful manual tuning of these choices, while our unified formulation automates this trade-off and is therefore more amenable to higher-order hierarchical models.
-6 Conclusion and Limitation
-We presented pragmatic curiosity, an AIF-based paradigm for hybrid learning-optimization under expensive black-box evaluations.
-By selecting actions via EFE minimization, the approach unifies goal-seeking and information-seeking within a single acquisition objective.
-Across constrained system identification, targeted active search, and composite optimization with unknown preferences, pragmatic curiosity consistently outperformed strong BO-type and BED-type baselines, improving estimation accuracy, critical-region coverage, and final solution quality under fixed budgets.
-The limitation of our approach stems from the specification of the problem-dependent energy/preference model; misspecification can bias acquisition and degrade performance.
-Performance also inherits assumptions of the underlying surrogate and observation models; strong model mismatch or non-stationarity can compromise both uncertainty quantification and preference-induced guidance.
-Future work includes extending the evaluation of this paradigm to multi-agent, long-horizon, or multi-fidelity settings.
-Impact Statement
-This work advances sample-efficient decision-making for engineering and scientific systems where evaluations are expensive and objectives are coupled to uncertainty.
-By unifying goal-seeking and information-seeking within a single EFE objective, pragmatic curiosity provides a general acquisition-design template that can reduce experimental cost, accelerate discovery, and improve reliability in workflows such as safe system identification, targeted failure-region discovery, and preference-driven design.
-In domains like environmental monitoring, autonomous-system testing, and energy-resource allocation, the practical benefit is fewer costly or risky evaluations for a given level of performance and confidence, which can translate into reduced resource consumption and safer experimentation.
-Potential negative impacts stem from misuse or mis-specification of the preference/energy model: poorly chosen goals or biased preferences can steer exploration toward harmful regions (e.g., unsafe operating conditions) or encode inequitable outcomes in preference-driven optimization.
-In addition, the ability to more efficiently identify system vulnerabilities could be misused in adversarial settings.
-These risks motivate careful deployment practices, including explicit safety constraints, transparent preference specification, robustness checks under model mismatch, and domain-appropriate oversight when the method is applied to safety-critical or socially sensitive applications.
-Appendix A Related Work
-Several independent attempts have been made to handle the hybrid objectives of goal seeking and learning.
-Below, we discuss some key prior works, originating from three viewpoints: (1) BO-based settings, (2) BED (also known as Bayesian active learning, BAL) based works, and (3) some works that have acknowledged the synergy and attempted to present a unifying viewpoint, using adaptive sampling and AIF-based approaches.
-On the BO side, Russo and Van Roy (2018) proposed information-directed sampling (IDS) to online optimization problems by minimizing the ratio between squared expected single-period regret and a measure of information gain.
-Hvarfner et al.
-(2023) introduced a statistical distance-based active learning (SAL) criterion into the BO loop to actively learn the Gaussian process hyperparameter even as it searches for the optimum.
-Tu et al.
-(2022); Hvarfner et al.
-(2022a) propose an information-theoretic acquisition function for multi-objective BO, which considers the joint information gain for the optimal set of inputs and outputs.
-Overall, these works acknowledge and demonstrate the value of information gain in enhancing the capabilities of goal-seeking strategies, with an ultimate focus on optimization.
-However, they do not explicitly comment on the synergy of goal seeking and learning, or formalize the acquisition strategies as a way to explicitly balance both objectives.
-On the BED (also known as Bayesian active learning, BAL) side, Smith et al.
-(2023) proposed an expected predictive information gain (EPIG) criterion that focuses on information gain in model predictions, mitigating classical BAL’s tendency to select out-of-distribution or low-relevance queries by accounting for an input data distribution.
-Lookman et al.
-(2019) discusses the performance of various BAL and BO strategies for material design application, noting their connection with adaptive sampling, and simultaneous exploration and exploitation.
-This shows that the realistic application of BO/BED-based approaches often necessitates a formalized viewpoint unifying each distinct objective.
-These efforts begin to show the growing synergy between BO and BED techniques, where information-gain criteria are leveraged to enhance optimization and vice versa.
-In search of a unifying perspective for these techniques, Di Fiore et al.
-(2023) views BO and BAL as symbiotic adaptive sampling techniques driven by common principles of ‘goal-driven learning’.
-However, their viewpoint is limited to BO and BED as separate realizations of adaptive sampling, and they do not discuss a way to combine both objectives in a single acquisition strategy.
-In this paper, we present a unified perspective that arises naturally from the principles of Active Inference (AIF) (Friston, 2010; Friston et al., 2017).
-Originally developed in computational neuroscience, AIF has been successfully applied in fields such as robotics (Lanillos et al., 2021), cognitive science (Smith et al., 2022), ecosystem modeling (Friston et al., 2024), and goal-directed planning in partially observable environments (Matsumoto et al., 2025).
-These works leverage the ability of AIF to accommodate goal-directed and epistemic considerations, which we utilize as the backbone for our unifying mechanism discussed in this paper.
-Appendix B Proof of Theorem 3.1
-Proof.
-Recall that
-$G=\underbrace{-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x)\right]}_{\text{Term 1}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y|x)}_{\text{Term 2}}$
-Since our purpose is to align the true outcome distribution given the decision $x$ (i.e., $p(y|x)$) with the prior preference distribution $p(y)$, we replace the second term of $G$ by $\mathbb{E}_{q(y|x)}\log p(y)$.
-It is commonly referred to as pragmatic value, which scores the anticipated surprisal of the future outcomes given a predictive distribution $q(y|x)$.
-For the first term of $G$, add and subtract $\log q(s|x,y)$ inside the expectation, we get
-$\begin{split}\text{Term 1}&=-\mathbb{E}_{q(y,s|x)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)+\log q(s|x,y)-\log q(s|x)\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\mathbb{E}_{q(s|x,y)}\left[\log p(s|x,y)-\log q(s|x,y)\right]+\mathbb{E}_{q(s|x,y)}\left[\log q(s|x,y)-\log q(s|x)\right]\right]\\ &=-\mathbb{E}_{q(y|x)}\left[\underbrace{-D_{\text{KL}}(q(s|x,y)||p(s|x,y))}_{\text{Term 3}}+\underbrace{D_{\text{KL}}(q(s|x,y)||q(s|x))}_{\text{Term 4}}\right].\end{split}$ (5)
-Since the surrogate model $q(s)$ usually will not be updated until a new outcome $y$ is observed, it doesn’t depend on the decision variable $x$ solely, i.e., $q(s|x)=q(s)$.
-Thus Term 4 becomes $D_{\text{KL}}(q(s|x,y)||q(s))$, which is commonly referred to as the epistemic value, or the expected information gain of a state when it is conditioned on expected observations.
-Term 3 is, in general, intractable since it depends on the true posterior distribution $p(s|x,y)$ that is often unknown.
-The best surrogate model we can have is $q(\cdot)=p(\cdot|\mathcal{D}_{t})$ that is constructed from all available data $\mathcal{D}_{t}$.
-Thus if we treat this surrogate model $q(s)$ as the best approximation of the true model $p(s)$, i.e., $q(s)\approx p(s)$, and the surrogate and true models both follow the Bayes’ rule, i.e., $q(s|x,y)=\frac{p(x,y|s)q(s)}{\int p(x,y|s)q(s)ds},\quad p(s|x,y)=\frac{p(x,y|s)p(s)}{\int p(x,y|s)p(s)ds},$ then Term 3 vanishes.
-∎
-Appendix C Proof of Lemma 3.2
+Fig.
+4 compares PraC with BOPE-style stage-wise variants on energy allocation.
+BOPE alternates between preference exploration and experimentation through manually specified stages or switching rules.
+PraC instead integrates outcome learning, preference learning, and preference-guided optimization at every iteration through a single acquisition.
+The comparison shows that stage-wise methods are sensitive to their switching schedules, while PraC more reliably discovers high-preference regions.
+Thus, the hierarchical version of PraC not only learns what the world does, but also learns what matters, without requiring a manually staged separation between learning and optimization.
+6 Conclusion
+This paper introduced Pragmatic Curiosity (PraC), an active-inference framework for hybrid learning and optimization.
+PraC addresses settings in which an agent cannot cleanly separate what it needs to learn from what it needs to do.
+By coupling epistemic information-seeking with a regret-based pragmatic potential, PraC provides a unified acquisition principle for deciding which information is worth acquiring because of how it may improve action.
+The framework exposes three operational design choices: symbols, the latent quantities whose clarification can change action; regrets, the potential landscapes that encode task value or shortfall; and curiosity, the exchange rate between information and pragmatic value.
+These choices provide a common language for interpreting existing BO and BED acquisition rules, while enabling new designs for hybrid problems where learning, optimization, and preference formation are coupled.
+We instantiated PraC across decision-oriented monitoring, targeted active search, and composite Bayesian optimization with unknown preferences.
+Across these regimes, PraC reduced downstream decision risk, improved coverage of critical outcome regions, and jointly learned predictive and preference structures without relying on task-specific staging rules.
+More broadly, PraC suggests that curiosity should not be treated merely as an exploration heuristic.
+When grounded in task-relevant symbols and regulated by regret-based potentials, curiosity becomes a mechanism for refining knowledge and action together.
+Future work includes non-myopic extensions, stronger theoretical guarantees, scalable approximations, and applications to embodied robotic systems.
+Appendix A Proof of Lemma 4.1 (Representation Compression)
 To prove Lemma 4.1, we first introduce a lemma:
-Lemma C.1.
+Lemma A.1.
 For any (finite or infinite) set $\mathcal{X}$, after a few new measurements $(\mathbf{X},\mathbf{Y}),\mathbf{X}\subseteq\mathcal{X}$ the KL divergence between $p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))$ and $p(f_{\mathcal{X}}|\mathcal{D})$ is
-$D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]=\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],$
+$D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]=\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],$
 where $L(\mathbf{Y}|f_{\mathbf{X}})$ is the likelihood of observations, $L(\mathbf{Y})=\int L(\mathbf{Y}|f_{\mathbf{X}})p(f_{\mathbf{X}}|\mathcal{D})df_{\mathbf{X}}$ the marginal likelihood.
@@
 The KL divergence between $p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))$ and $p(f_{\mathcal{X}}|\mathcal{D})$ can be given as
-$\begin{split}&D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]\\ =&D_{\text{KL}}[p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})]\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\int\frac{p(f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (6)
+$\begin{split}&D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]\\ =&D_{\text{KL}}[p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})]\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathcal{X}\setminus\mathbf{X}},f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathcal{X}\setminus\mathbf{X}}df_{\mathbf{X}}\\ =&\int p(f_{\mathbf{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\int\frac{p(f_{\mathbf{X}}|\mathcal{D})L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}df_{\mathbf{X}}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (6)
 where $L(\mathbf{Y}|f_{\mathbf{X}})$ is the likelihood of observations, $L(\mathbf{Y})=\int L(\mathbf{Y}|f_{\mathbf{X}})p(f_{\mathbf{X}}|\mathcal{D})df_{\mathbf{X}}$ the marginal likelihood.
@@
 Thus, we need to resort to a more general definition for KL divergence based on the Radon-Nikodym derivative:
-Definition C.2.
+Definition A.2.
 If $P$ and $Q$ are probability measures over a set $\mathcal{X}$, and $P$ is absolutely continuous with respect to $Q$, then the KL divergence from $P$ to $Q$ is defined as
-$D_{\text{KL}}[P||Q]=\int_{\mathcal{X}}\log(\frac{dP}{dQ})dP,$
+$D_{\text{KL}}[P\|Q]=\int_{\mathcal{X}}\log(\frac{dP}{dQ})dP,$
 where $\frac{dP}{dQ}$ is the Radon–Nikodym derivative of $P$ with respect to $Q$, and provided the expression on the right-hand side exists.
@@
 Now the KL divergence between $\hat{P}$ and $P$ is quantified as
-$\begin{split}&D_{\text{KL}}[P(f_{\mathcal{X}})||\hat{P}(f_{\mathcal{X}})]\\ =&\int_{f_{\mathcal{X}}}\log(\frac{dP}{d\hat{P}}(f_{\mathcal{X}}))dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathcal{X}})\\ =&\int_{f_{\mathbf{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathbf{X}})\\ =&\mathbb{E}_{\hat{P}(f_{\mathbf{X}})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (7)
-which has the exact same form as (6).
+$\begin{split}&D_{\text{KL}}[P(f_{\mathcal{X}})\|\hat{P}(f_{\mathcal{X}})]\\ =&\int_{f_{\mathcal{X}}}\log(\frac{dP}{d\hat{P}}(f_{\mathcal{X}}))dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})dP(f_{\mathcal{X}})\\ =&\int_{f_{\mathcal{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathcal{X}})\\ =&\int_{f_{\mathbf{X}}}\log(\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})})\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\hat{P}(f_{\mathbf{X}})\\ =&\mathbb{E}_{\hat{P}(f_{\mathbf{X}})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}],\end{split}$ (7)
+which has the exact same form as equation 6.
 Therefore, we can conclude that regardless of the set $\mathcal{X}$ being finite or infinite, the KL divergence between the prior and posterior only depends on the evaluations of the observed data.
@@
 Proof.
-According to Lemma C.1,
-$\begin{split}I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D})=&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}[D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))||p(f_{\mathcal{X}}|\mathcal{D})]]\\ =&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{L(\mathbf{Y})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\mathbb{E}_{L(\mathbf{Y})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]L(\mathbf{Y})d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int L(\mathbf{Y}|f_{\mathbf{X}})\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[D_{\text{KL}}[L(\mathbf{Y}|f_{\mathbf{X}})||L(\mathbf{Y})]]\\ =&I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}).\end{split}$ (8)
-Appendix D Experimental Details
+According to Lemma A.1,
+$\begin{split}I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D})=&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}[D_{\text{KL}}[p(f_{\mathcal{X}}|\mathcal{D}\cup(\mathbf{X},\mathbf{Y}))\|p(f_{\mathcal{X}}|\mathcal{D})]]\\ =&\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},\mathcal{D})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{L(\mathbf{Y})}\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\mathbb{E}_{L(\mathbf{Y})}[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int[\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}]L(\mathbf{Y})d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}\int L(\mathbf{Y}|f_{\mathbf{X}})\log\frac{L(\mathbf{Y}|f_{\mathbf{X}})}{L(\mathbf{Y})}d\mathbf{Y}\\ =&\mathbb{E}_{p(f_{\mathbf{X}}|\mathcal{D})}[D_{\text{KL}}[L(\mathbf{Y}|f_{\mathbf{X}})\|L(\mathbf{Y})]]\\ =&I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}).\end{split}$ (8)
+Appendix B Re-interpretation of GP-UCB
+The reinterpretations of most of the acquisition strategies in Table 1 are straightforward according to their definitions.
+However, placing a rather intuitive GP-UCB strategy within this framework seems implicit.
+To reveal their connection, we resort to Lemma 4.1.
+Then if we assume constant Gaussian noises $\mathcal{N}\sim(0,\sigma^{2})$ for the observations, we have
+$\begin{split}I(f_{x};y|\mathcal{D}_{t})&=h(y\mid\mathcal{D}_{t})-H(y|f_{x},\mathcal{D}_{t})\\ &=\frac{1}{2}\log(1+\sigma^{-2}\sigma_{t}^{2}(x)),\end{split}$
+where $\sigma_{t}^{2}(x)$ is the variance evaluated on the GP model $p(f_{x}|\mathcal{D}_{t})$.
+When further assuming that the GP kernel $\kappa_{t}(x,x^{\prime})\leq 1,\forall x,x^{\prime}\in\mathcal{X}$, then $0\leq\sigma_{t}^{2}(x)\leq\kappa_{t}(x,x^{\prime})\leq 1$, which gives
+$\log(1+\sigma^{-2}\sigma_{t}^{2}(x))\geq\log(1+\sigma^{-2})\sigma_{t}^{2}(x).$
+If we choose $\beta=\frac{1}{2}\log(1+\sigma^{-2})$, then the epistemic term in PraC, i.e., $I(f_{x};y|\mathcal{D}_{t})$, provides an upper bound of the square of the exploration term $\beta^{1/2}\sigma_{t}(x)$ in GP-UCB.
+This reveals the close relationship between GP-UCB and AIF, showing that the uncertainty bonus in GP-UCB can be viewed as a first-order surrogate for an information-theoretic epistemic term.
+Appendix C Experimental Details
 This appendix provides a comprehensive overview of the simulation environment, model parameters, and hyper-parameter choices used to generate the results in this paper.
 The experiments were designed to be reproducible given the configurations outlined below.
-D.1 Simulation Environment
-All simulations for the perception failure evaluation in CARLA (Section D.3) were conducted on a Linux workstation with Ubuntu 22.04 LTS equipped with an Intel 13th Gen Core i7-13700KF CPU (16 cores, 24 threads, up to 5.4 GHz) and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM).
+C.1 Simulation Environment
+All simulations for the perception failure evaluation in CARLA (Section C.3) were conducted on a Linux workstation with Ubuntu 22.04 LTS equipped with an Intel 13th Gen Core i7-13700KF CPU (16 cores, 24 threads, up to 5.4 GHz) and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM).
 The system ran CARLA simulations using CUDA 12.2 and NVIDIA driver version 535.230.02.
@@
 The Monte Carlo samples are drawn from the posteriors for each model to approximate the expectations of acquisition functions.
-D.2 Constrained System Identification
-D.2.1 Plume Field Model and Parameters
+C.2 Decision-Oriented Plume Monitoring
+This appendix provides implementation details and additional analysis for the decision-oriented plume-monitoring experiments in Section 5.1.
+All three tasks share the same sequential decision structure.
+The agent maintains a posterior distribution over latent environmental hypotheses $\theta$, selects one sensor query $x$ at each iteration, observes a plume count $y$, updates the posterior, and evaluates the downstream decision induced by the updated belief.
+C.2.1 Plume Field Model
 We consider the monitoring of a chemical plume field, where multiple plume sources generate plume particles that can be measured by sensors.
@@
 where $\Delta t$ is the time taken to obtain a measurement.
-General Field and Sensor Parameters:
-  • Grid Size: 100 × 100 units.
-  • Sensor Size ($a$): 1.0 units.
-  • Measurement Time ($\Delta t$): 1.0 seconds.
-Source Parameters: Eight potential plume sources were defined for the experiments. Their specific physical parameters are detailed as:
-Source $\theta$ $R_{s}$ $\gamma$ $V$ $D$
-1 [20.0, 20.0] 100.0 50.0 [0.5, 0.5] 10
-2 [30.0, 80.0] 100.0 60.0 [-0.3, 0.2] 15
-3 [45.0, 55.0] 15.0 50.0 [0.5, 0.5] 10
-4 [50.0, 50.0] 18.0 30.0 [-0.3, 0.2] 15
-5 [55.0, 45.0] 16.0 40.0 [0.2, -0.4] 12
-6 [48.0, 52.0] 17.0 35.0 [0.1, 0.1] 11
-7 [52.0, 55.0] 14.0 45.0 [-0.1, -0.1] 13
-8 [52.0, 52.0] 18.0 40.0 [0.1, -0.1] 13
-D.2.2 Task-Specific Configurations
-Source Localization.
-  • Goal: Estimate the unknown location $\theta=[x,y]$ of a single active source.
-  • Ground Truth: Source 1 was configured as the single active source.
-  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
-  • Hypothesis Space: A discrete grid of potential locations spanning [0,100]×[0,100] with a resolution of 5 units, resulting in a 20 × 20 grid of 400 hypotheses.
-  • Degree of curiosity: $\beta=0.5$.
-Wind Estimation.
-  • Goal: Estimate the unknown wind vector $V=[v_{x},v_{y}]$ for a single source with a known location.
-  • Ground Truth: Source 2 was used, with its true wind vector set to [-0.3, 0.2].
-  • Maximum Sensor Threshold ($y_{max}$): 60.0 hits/second.
-  • Hypothesis Space: A discrete grid of potential wind vectors spanning [-1,1]×[-1,1] with a resolution of 0.1 units, resulting in a 20 × 20 grid of 400 hypotheses.
-  • Degree of curiosity: $\beta=1.0$.
-Active Source Identification.
-  • Goal: Identify the subset of active sources from a set of six potential sources with known locations.
-  • Ground Truth: Sources 3-8 were used, with Sources 3, 5, 6, and 8 were set as active.
-  • Maximum Sensor Threshold ($y_{max}$): 30.0 hits/second.
-  • Hypothesis Space: The set of all possible on/off combinations for the six sources. This is a discrete space with $2^{6}=64$ unique hypotheses.
-  • Degree of curiosity: $\beta=5.0$.
-D.3 Constrained Active Search
-D.3.1 Perception in self driving simulation CARLA
+C.2.2 Task 1: Source Response Localization
+The first task asks the agent to localize a plume source well enough to dispatch a response team near it.
+The true source is located at $(35,65)$.
+The latent hypotheses $\theta$ consist of source-location hypotheses on a grid over the $100\times 100$ domain with spacing $5$, together with source-strength multipliers
+$[0.4,0.7,1.0,1.6].$
+The response actions form a coarser grid with spacing $20$.
+The prior is intentionally multimodal, with modes near $(75,25)$ and $(35,65)$, so the agent must resolve a decision-relevant ambiguity rather than simply exploit a unimodal belief.
+For a response action $a$ and hypothesis $\theta$, let $\ell_{\theta}$ denote the source location specified by $\theta$. The downstream loss is the normalized squared response distance,
+$L_{\mathrm{loc}}(a,\theta)=\frac{\|a-\ell_{\theta}\|_{2}^{2}}{100^{2}+100^{2}}.$
+Given posterior $q_{t}(\theta)=p(\theta\mid\mathcal{D}_{t})$, the Bayes risk is
+$BR(p_{t})=\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{loc}}(a,\theta),$
+and the Bayes action is
+$a_{t}^{\star}=\arg\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{loc}}(a,\theta).$
+For dynamic curiosity schedule, the default dynamic settings are
+$q=0.75,\quad k_{\beta}=1,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=10^{-3},\quad\beta_{\max}=10,\quad U^{\star}=0.$
+We report:
+  • Bayes risk: $BR(p_{t})$.
+  • Response loss: realized loss of the Bayes action under the true source,
+    $L_{\mathrm{loc}}(a_{t}^{\star},\theta_{\mathrm{true}}).$
+  • Response success: whether the selected response action lies within one response-grid spacing of the true source,
+    $\mathbf{1}\{\|a_{t}^{\star}-\ell_{\theta_{\mathrm{true}}}\|_{2}\leq 20\}.$
+  • Parameter error: Euclidean error between the posterior MAP source location and the true source location.
+C.2.3 Task 2: Consequence-Weighted Dispatch
+The second task uses the same response-decision structure, but errors in high-consequence regions are more costly.
+The true source is located at $(68,72)$.
+The prior has modes near $(34,66)$, $(68,72)$, and $(76,24)$, and the high-consequence region is centered near the true source.
+Let $C(\ell_{\theta})$ denote the consequence weight associated with the source location under hypothesis $\theta$.
+In the implementation, this weight is defined by a baseline term plus Gaussian consequence regions,
+$C(\ell_{\theta})=C_{0}+\sum_{j}w_{j}\exp\left(-\frac{\|\ell_{\theta}-c_{j}\|_{2}^{2}}{2\sigma_{j}^{2}}\right),$
+with an additional source-strength multiplier. The consequence-weighted downstream loss is
+$L_{\mathrm{cw}}(a,\theta)=C(\ell_{\theta})\frac{\|a-\ell_{\theta}\|_{2}^{2}}{100^{2}+100^{2}}.$
+Given posterior $q_{t}(\theta)$, the Bayes risk is
+$BR(p_{t})=\min_{a}\sum_{\theta}q_{t}(\theta)L_{\mathrm{cw}}(a,\theta).$
+For dynamic curiosity schedule, the default dynamic settings are
+$q=0.75,\quad k_{\beta}=1,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=10^{-3},\quad\beta_{\max}=10,\quad U^{\star}=0.$
+We report:
+  • Bayes risk: posterior expected consequence-weighted loss of the Bayes action.
+  • Response loss: realized consequence-weighted response loss under the true source.
+  • Response success: the same geometric success criterion as in source response localization.
+  • Parameter error: Euclidean MAP source-location error.
+C.2.4 Task 3: Active Source Prioritization
+The third task models limited repair or inspection resources.
+There are six possible sources, and the true active sources are indices $[4,5]$.
+The repair budget is $k=2$.
+The source consequence weights are
+$[70,65,5,5,180,170],$
+and the prior activity probabilities are
+$[0.82,0.80,0.40,0.40,0.15,0.15].$
+This creates a decision tension: the prior favors high-probability decoys, while the true active sources have high consequence weights but low prior activity.
+The posterior is over nonempty active-source subsets $\theta$. Let
+$z_{i}(\theta)\in\{0,1\}$
+indicate whether source $i$ is active under hypothesis $\theta$. The posterior marginal activity probability of source $i$ is
+$p_{t,i}=\sum_{\theta}q_{t}(\theta)z_{i}(\theta).$
+With zero false-positive cost, the expected missed-source risk before selecting repairs is
+$\sum_{i}w_{i}p_{t,i}.$
+If the selected repair set is $S$, its posterior expected benefit is
+$\sum_{i\in S}w_{i}p_{t,i}.$
+The Bayes action selects the top $k$ sources by $w_{i}p_{t,i}$, and the Bayes risk is
+$BR(p_{t})=\sum_{i}w_{i}p_{t,i}-\max_{|S|=k}\sum_{i\in S}w_{i}p_{t,i}.$
+For dynamic curiosity schedule, the default dynamic settings are
+$q=0.9,\quad k_{\beta}=2,\quad\varepsilon_{I}=10^{-8},\quad\varepsilon_{U}=10^{-8},\quad\beta_{\min}=1,\quad\beta_{\max}=10,\quad U^{\star}=0.$
+We report:
+  • Bayes risk: posterior expected weighted risk left unrepaired after selecting $k$ sources.
+  • Missed-source risk: realized total weight of true active sources not selected,
+    $\sum_{i\in A_{\mathrm{true}}\setminus S_{t}}w_{i}.$
+  • Top-$k$ recall:
+    $\frac{|S_{t}\cap A_{\mathrm{true}}|}{\min(k,|A_{\mathrm{true}}|)}.$
+  • Weighted top-$k$ recall:
+    $\frac{\sum_{i\in S_{t}\cap A_{\mathrm{true}}}w_{i}}{\sum_{i\in A_{\mathrm{true}}}w_{i}}.$
+  • Parameter error: symmetric-difference size between the MAP active-source set and the true active-source set.
+C.2.5 Baselines
+We compare the following methods.
+Random.
+Random selects a sensor query uniformly or according to the predefined random sampling protocol over the candidate set.
+Pure information gain.
+EIG selects
+$x_{t}=\arg\max_{x\in\mathcal{X}_{t}^{\mathrm{cand}}}\widehat{I}_{t}(x),$
+and therefore learns the latent environment without regard to downstream decision consequence.
+Decision-greedy.
+Greedy selects
+$x_{t}=\arg\max_{x\in\mathcal{X}_{t}^{\mathrm{cand}}}\widehat{\Delta BR}_{t}(x),$
+and therefore maximizes immediate expected downstream Bayes-risk reduction without explicitly valuing epistemic clarification.
+Fixed-$\beta$ PraC.
+Fixed PraC uses
+$\alpha_{t}(x)=\widehat{\Delta BR}_{t}(x)+5\,\widehat{I}_{t}(x).$
+Dynamic-$\beta$ PraC.
+Dynamic PraC uses the same acquisition but computes $\beta_{t}$ adaptively from the feedforward–feedback schedule described above.
+C.2.6 Analysis of Decision-Oriented Monitoring Results
+The decision-oriented plume-monitoring experiments are designed to test whether PraC learns environmental distinctions that matter for downstream action, rather than merely reducing parameter uncertainty.
+The detailed results are reported in Table 2.
+Across the three tasks, the results show three complementary behaviors: dynamic PraC improves downstream Bayes-risk reduction in the response-localization and source-prioritization tasks; decision-greedy can be competitive when posterior risk is already well aligned with the downstream loss; and pure information gain can fail when uncertainty reduction is not sufficiently targeted toward the decision-relevant ambiguity.
+Source response localization.
+In source response localization, dynamic PraC achieves the lowest final Bayes risk, $0.0025\pm 1.59\times 10^{-12}$, and also obtains zero parameter error.
+All non-random adaptive methods achieve the best realized response loss and response success, indicating that the response decision itself is relatively easy once the posterior identifies a response-equivalent source region.
+However, the Bayes-risk and parameter-error results reveal a sharper distinction among methods.
+Dynamic PraC resolves the latent ambiguity most completely, whereas fixed PraC and EIG retain residual MAP source-location error, and Greedy retains larger parameter error.
+This suggests that, in this task, downstream decision improvement and epistemic clarification are aligned, but dynamic curiosity still improves the posterior quality beyond what is required for merely selecting a successful response action.
+Consequence-weighted dispatch.
+The consequence-weighted dispatch task is more diagnostic because the posterior Bayes risk and the realized response loss no longer rank methods identically.
+Greedy achieves the lowest final Bayes risk, $0.0254\pm 0.00538$, because it directly optimizes the one-step expected reduction in posterior risk.
+However, dynamic PraC achieves the best realized response loss, $0.0347\pm 0.00771$, and the lowest parameter error, $7.62\pm 5.56$.
+This gap is important: posterior Bayes risk measures expected loss under the agent’s current belief, whereas realized response loss evaluates the downstream decision under the true latent hypothesis.
+A purely greedy policy can therefore reduce posterior risk quickly while still failing to collect information that would improve the realized decision under the true source.
+Dynamic PraC sometimes sacrifices immediate posterior-risk reduction to clarify the environmental hypothesis, which improves the eventual response decision in the realized environment.
+Active source prioritization.
+Active source prioritization creates a different form of decision ambiguity.
+The prior assigns high activity probabilities to some lower-consequence decoy sources, while the true active sources have lower prior probability but much larger consequence weights.
+Dynamic PraC achieves the lowest final Bayes risk, $2.86\pm 2.72$, improving over fixed PraC, Greedy, EIG, and Random.
+Dynamic and fixed PraC tie on missed-source risk, top-$k$ recall, and weighted top-$k$ recall, showing that both PraC variants reliably identify the high-consequence sources needed for the repair decision.
+By contrast, EIG and Random perform substantially worse on missed-source risk and weighted recall.
+This supports the central point of the decision-oriented formulation: information is useful only when it helps resolve the ambiguity that affects the downstream repair set.
+Pure information gain may reduce uncertainty over the active-source subset, but it does not necessarily prioritize distinctions that change the top-$k$ consequence-weighted decision.
+Comparison with pure information gain.
+EIG is a strong baseline when learning the latent hypothesis is well aligned with the downstream decision.
+This explains why it performs competitively in source response localization and matches fixed PraC on several metrics.
+However, EIG lacks an explicit mechanism for distinguishing decision-relevant uncertainty from decision-irrelevant uncertainty.
+In consequence-weighted dispatch and active source prioritization, this limitation becomes more visible: EIG can collect information about the environment without adequately targeting the distinctions that reduce consequence-weighted response loss or missed-source risk.
+PraC addresses this by coupling information gain with expected Bayes-risk reduction.
+Comparison with decision-greedy.
+The decision-greedy baseline isolates the pragmatic term by choosing queries that maximize $\widehat{\Delta BR}_{t}(x)$ without any explicit epistemic value.
+Its strong Bayes-risk performance in consequence-weighted dispatch shows that myopic posterior-risk reduction can be effective when the current belief already points toward useful response decisions.
+However, Greedy is more vulnerable when the current posterior is brittle or biased by the prior.
+In active source prioritization, Greedy is substantially worse than dynamic PraC in final Bayes risk, even though it remains competitive on some top-$k$ metrics.
+This indicates that myopic decision improvement may not be sufficient when the downstream action depends on resolving low-prior but high-consequence hypotheses.
+Why Bayes risk and realized loss can disagree.
+The distinction between Bayes risk and realized downstream loss is essential for interpreting these experiments.
+Bayes risk is the expected loss under the agent’s posterior belief, while realized loss is evaluated under the true latent hypothesis.
+A method can reduce Bayes risk by becoming confident under its current posterior, even if the posterior remains biased in a way that hurts realized performance.
+This is why Greedy can obtain the lowest final Bayes risk in consequence-weighted dispatch while dynamic PraC obtains the best realized response loss and parameter error.
+PraC’s epistemic term helps correct the posterior before committing too strongly to the current decision surrogate.
+Overall, the three tasks show that PraC is most beneficial when downstream action depends on resolving specific decision-relevant ambiguity.
+When the pragmatic objective and epistemic objective are aligned, PraC remains competitive with EIG and Greedy.
+When they diverge, dynamic PraC provides a mechanism for deciding when information is worth acquiring because it improves downstream action.
+This supports the role of PraC as a hybrid learning-and-optimization framework: the agent does not learn the environment for its own sake, nor does it merely optimize the current posterior decision; it learns the environmental distinctions that make better downstream action possible.
+Table 2: Full final performance metrics for decision-oriented plume monitoring over 20 seeds.
+Lower is better for Bayes risk, response loss, parameter error, and missed-source risk.
+Higher is better for response success, top-$k$ recall, and weighted top-$k$ recall.
+Bold indicates the best or tied-best mean within each task and metric.
+Dyn.
+PraC denotes dynamic-$\beta_{t}$ PraC; Fixed PraC denotes fixed-$\beta$ PraC with $\beta=5$.
+Source Response Localization
+Metric Dyn. PraC Fixed PraC Greedy EIG Random
+Bayes risk $\downarrow$ $\mathbf{0.0025\pm 1.59{\times}10^{-12}}$ $00377\pm 0.00254 0.0$$0463\pm 0.000765 0.0$$0377\pm 0.00254 0.0$$0124\pm 0.00495$
+Response loss $\downarrow$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $\mathbf{0.0025\pm 0}$ $0.0125\pm 0.00632$
+Response success $\uparrow$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $0.80\pm 0.40$
+Parameter error $\downarrow$ $\mathbf{0.00\pm 0}$ $2.00\pm 4.00$ $5.24\pm 5.25$ $2.00\pm 4.00$ $8.61\pm 5.09$
+Consequence-Weighted Dispatch
+Metric Dyn. PraC Fixed PraC Greedy EIG Random
+[…]
+Response loss $\downarrow$ $\mathbf{0.0347\pm 0.00771}$ $0424\pm 0.00945 0.0$$424\pm 0.00945 0.0$$424\pm 0.00945 0.0$$0463\pm 0.0144$
+Response success $\uparrow$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$ $\mathbf{1.00\pm 0}$
+Parameter error $\downarrow$ $\mathbf{7.62\pm 5.56}$ $8.86\pm 6.60$ $04\pm 3.06 8.8$$86\pm 6.60$ $17.8\pm 9.72$
+Active Source Prioritization
+Metric Dyn. PraC Fixed PraC Greedy EIG Random
+Bayes risk $\downarrow$ $\mathbf{2.86\pm 2.72}$ $3.59\pm 2.70$ $14.0\pm 10.6$ $23.3\pm 19.1$ $34.5\pm 17.0$
+Missed-source risk $\downarrow$ $\mathbf{34.0\pm 68.0}$ $\mathbf{34.0\pm 68.0}$ $36.0\pm 72.0$ $68.0\pm 83.3$ $316\pm 68.0$
+Top-$k$ recall $\uparrow$ $\mathbf{0.90\pm 0.20}$ $\mathbf{0.90\pm 0.20}$ $\mathbf{0.90\pm 0.20}$ $0.80\pm 0.245$ $0.10\pm 0.20$
+Weighted top-$k$ recall $\uparrow$ $\mathbf{0.903\pm 0.194}$ $\mathbf{0.903\pm 0.194}$ $0.897\pm 0.206$ $806\pm 0.238 0.0$$0971\pm 0.194$
+C.2.7 Diagnostics of Dynamic Curiosity
+In addition to the main performance curves, we report scheduler diagnostics to verify that dynamic curiosity behaves as intended. The diagnostic plots include:
+  • $\beta_{t}$, the scheduled curiosity coefficient;
+  • $\beta_{t}^{\mathrm{ff}}$, the feedforward exchange-rate scale;
+  • $\beta_{t}^{\mathrm{fb}}$, the feedback activation factor;
+  • $\beta_{t}\widehat{I}_{t}(x_{t})$, the selected effective epistemic pressure.
+Figure 5: Diagnostics of dynamic curiosity scheduler. Shaded regions show mean $\pm 1$ standard error over 20 seeds.
+These diagnostics distinguish three regimes.
+When decision-relevant uncertainty is high, the feedback activation keeps curiosity active.
+When useful information is costly relative to downstream Bayes-risk reduction, the feedforward scale increases.
+When the decision-symbol posterior concentrates, the feedback activation suppresses curiosity even if raw information-gain ratios become numerically large.
+This prevents late-stage ratio explosions and makes the dynamic schedule interpretable.
+Role of dynamic curiosity.
+The dynamic schedule gives $\beta_{t}$ an operational meaning as an exchange rate between epistemic value and downstream Bayes-risk reduction.
+The feedforward scale estimates how large curiosity must be for information gain to compete with pragmatic improvement, while the feedback activation suppresses curiosity as decision-relevant uncertainty decreases.
+This prevents $\beta_{t}$ from acting as a fixed exploration knob.
+Instead, the relevant quantity is the effective epistemic pressure $\beta_{t}\widehat{I}_{t}(x)$.
+The empirical results are consistent with this interpretation: dynamic PraC improves posterior quality in source response localization, improves realized downstream loss in consequence-weighted dispatch, and achieves the lowest Bayes risk in active source prioritization.
+C.3 Targeted Active Search
+C.3.1 Perception in self driving simulation CARLA
 We consider the failure discovery for YOLO-based object detection (Jiang et al., 2022; Redmon and Farhadi, 2018) in the CARLA simulator (Dosovitskiy et al., 2017).
@@
 6 shows examples of the discussed object detection failures we aim to discover.
-Figure 4: Examples of missed object detection by YOLO due to two reasons considered in perception failure case study in Section D.3.
+Figure 6: Examples of missed object detection by YOLO due to two reasons considered in perception failure case study in Section C.3.
 Fig (left to right): example of Failure-1 (distance), Failure-2 (poor light) and Failure-1 and Failure-2 both in one scene (distance and poor light), respectively.
@@
 Each scene has two cars and a pedestrian.
-D.3.2 LLM-based evaluation for CARLA
+C.3.2 LLM-based evaluation for CARLA
 We use LLM to perform evaluations for whether a generated scene corresponds to a failure due to specific type.
@@
 The output of the LLM is used to obtain a binary number for each scene which is composed to give a scenario specific cost function $c_{1},c_{2}$.
-D.3.3 Task-Specific Configurations
+C.3.3 Task-Specific Configurations
   • Goal: We consider two cost functions $c_{1},c_{2}$ associated with each type of failure.
@@
 The list of objects detected and reason for incomplete detection for the image are as follows: Objects detected: {objects} Reason: {reason}
-D.4 Composite Bayesian Optimization
-D.4.1 Preference Evaluation
+C.4 Composite Bayesian Optimization
+C.4.1 Preference Evaluation
 We simulate human-in-the-loop or policy-driven decision-making via pairwise preference queries.
@@
 An initial set of 1 pairwise preferences is randomly sampled to initialize the model. Each step of the optimization selects new pairs to query, guided by the used acquisition strategy.
-D.4.2 Task-Specific Configurations
+C.4.2 Task-Specific Configurations
 Vehicle Safety.
@@
   • Testbed: See Tanabe and Ishibuchi (2020) for details.
-  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[1864.7202,11.8199,0.2904]$.
+  • Ground Truth: $g(y)=(y-y^{\ast})^{2}$, where $y^{\ast}=[1864.7202,11.8199,0.2904]$.
   • Degree of curiosity: $\beta=\gamma=1.0$.
@@
   • Testbed: See Liang and Lai (2021) for details.
-  • Ground Truth Preference: $g(y)=-(y-y^{\ast})^{2}$, where $y^{\ast}=[25.935,57.612,935.5]$.
+  • Ground Truth: $g(y)=(y-y^{\ast})^{2}$, where $y^{\ast}=[25.935,57.612,935.5]$.
   • Degree of curiosity: $\beta=\gamma=1.0$.
@@
   • Testbed: IEEE 30-bus network in pandapower library.
-  • Ground Truth Preference: $g(y)=a^{\intercal}y$, where a=[1,1,2,1]a=[1,-1,2,1]
+  • Ground Truth Preference: $g(y)=a^{\intercal}y$, where a=[1,1,2,1]a=[-1,1,-2,-1]
   • Degree of curiosity: $\beta=\gamma=1.0$.
@@
 Resilience Assesses the percentage of time that all bus voltages remain within safe operating limits under perturbations (e.g., load uncertainty or line outages).
-D.4.3 Extended Baseline Comparison with BOPE
+C.4.3 Comparison with BOPE
 To highlight the benefits of jointly learning and optimizing, rather than separating these into stages, we extend the baseline comparison with BOPE from Lin et al. (2022).
@@
   • BOPE-IV: A BOPE variant that uses standard preference exploration (qEUBO) and selects experiments exclusively with qNEI, fully refitting both outcome and preference GPs after each update.
-Figure 5: Extended baseline comparison with BOPE for energy resource allocation. Error bars represent $\pm 1$ standard deviation over 5 seeds.
+Figure 7: Extended baseline comparison with BOPE for energy allocation. Error bars represent $\pm 1$ standard deviation over 20 seeds.
 Figure 7 reports their preference scores over random sampling (RS).
-It is evident that our active-inference method for hierarchical GPs (HGP) consistently discovers higher-preference regions after a brief initial exploration phase, while BOPE variants are highly sensitive to their stage-wise design choices.
+It is evident that our method (PraC) consistently discovers higher-preference regions after a brief initial exploration phase, while BOPE variants are highly sensitive to their stage-wise design choices.
 BOPE-I, being preference-driven in the first phase, initially attains higher preference values but fails to balance exploration and exploitation, leading to poor convergence in the second half; its performance is also sensitive to the precise switching point.
 BOPE-II explores first and then optimizes, achieving better final performance than BOPE-I, but the strict separation between exploration and exploitation still yields suboptimal outcomes.
-BOPE-III mixes both aspects but remains more exploitation-centric, performing better than BOPE-I/II yet still below HGP.
-BOPE-IV and HGP share the idea of refitting both models at each step, but BOPE-IV converges to a lower-preference solution.
+BOPE-III mixes both aspects but remains more exploitation-centric, performing better than BOPE-I/II yet still below PraC.
+BOPE-IV and PraC share the idea of refitting both models at each step, but BOPE-IV converges to a lower-preference solution.
 In contrast, our acquisition strategy jointly leverages information from both the outcome and preference models at every iteration, leading to higher sample efficiency and more reliable discovery of high-preference regions.
-Appendix E Use of Large Language Models
-We use LLM to revise and polish drafts of the introduction, abstract, conclusion, and limitations sections to enhance clarity, flow, and accessibility for a broader audience.
-We also use LLM as an evaluator for the experiments of failure discovery in CARLA, which is detailed in Appendix D.3.